### 11 Mar 2026 v1

# On Utility Maximization under Multivariate Fake Stationary Affine Volterra Models

Abstract
This paper is concerned with Merton’s portfolio optimization problem in a Volterra stochastic environment described by a multivariate fake stationary Volterra–Heston model. Due to the non-Markovianity and non-semimartingality of the underlying processes, the classical stochastic control approach cannot be directly applied in this setting. Instead, the problem is tackled using a stochastic factor solution to a Riccati backward stochastic differential equation (BSDE). Our approach is inspired by the martingale optimality principle combined with a suitable verification argument. The resulting optimal strategies for Merton’s problems are derived in semi-closed form depending on the solutions to time-dependent multivariate Riccati-Volterra equations. Numerical results on a two dimensional fake stationary rough Heston model illustrate the impact of stationary rough volatilities on the optimal Merton strategies.
Keywords: Affine Volterra Processes, Stochastic Control, Martingale Optimality Principle, Backward Stochastic Differential Equations (BSDE), Fractional Differential Equations, Riccati Equations, Functional Integral Equation.
Mathematics Subject Classification (2020): 34A08, 34A34, 45D05, 60G10, 60G22, 60H10, 91B70, 91G80,93E20
1 Introduction
The modeling of asset price dynamics has undergone a paradigm shift with the empirical observation that both implied and realized volatilities of major financial indices exhibit significantly rougher sample paths (Gatheral et al., 2018) than those generated by classical Brownian-motion-based models. This observation has sparked a rapidly expanding body of research on rough volatility models. Motivated by the widespread practical success of the celebrated Heston (1993) stochastic volatility framework, several rough extensions have been proposed, most notably the rough Heston model (El Euch and Rosenbaum, 2019) builds on market microstructure and its generalisation, the Volterra Heston model in Abi Jaber et al. (2019). Recent remarkable advances include the introduction of the so-called fake stationary Volterra Heston models in Gnabeyeu et al. (a, b) with the aim of providing a unified and consistent framework that captures both short- and long-maturity behaviors and enables robust fitting across the entire term structure. This broader class of volatility models, encompassing the aforementioned specifications, is obtained by modeling the volatility process as a stochastic Volterra equation of convolution type with a time-dependent diffusion coefficient. Therefore, this paper focuses on the financial market within the fake stationary affine Volterra model.
Substantial progress has recently been achieved in the study of option pricing problems and asymptotic analysis under rough volatility dynamics. By contrast, portfolio optimization in such models remains comparatively underexplored, although it has attracted growing interest in recent years. Notable contributions include Fouque and Hu (2019); Bäuerle and Desmettre (2020); Han and Wong (2020), which investigate optimal investment problems with power utility in fractional Heston-type models. Despite these advances, the overwhelming majority of developments in rough volatility, whether for asset modeling, derivative pricing, or portfolio selection, have been confined to the mono-asset case. From a practical standpoint, however, multi-asset allocation with correlated sources of risk constitutes a fundamental dimension of portfolio management; see, for instance, Buraschi et al. (2010).
Merton’s portfolio optimization problem which consists in maximizing an investor’s expected utility from terminal wealth with respect to a given utility function has served as a cornerstone in mathematical finance. It is the most classic financial economic approach to understand how the market volatility affect investment demands. In the classical Heston (1993) stochastic volatility framework, this problem was solved explicitly in Kraft (2005), building on the representation results of Zariphopoulou (2001), and extensions to general affine stochastic volatility models were obtained in Kallsen and Muhle-Karbe (2010) using stochastic control theory. In the Volterra framework, however, the volatility process is non-Markovian, which prevents the direct application of the classical stochastic control methods based on the Hamilton–Jacobi–Bellman (HJB) partial differential equation. In order to circumvent this difficulty, Han and Wong (2020), inspired by Fouque and Hu (2019), adopt a martingale distortion ansatz and apply the martingale optimality principle to derive explicit optimal investment strategies in a single-asset or mono-asset Volterra–Heston model.
Motivated by several important empirical stylized facts about real financial markets such as choice among multiple assets, rough volatility behavior, correlations across stocks or assets and leverage effects (i.e., correlation between a stock and its volatility), multivariate rough volatility models have recently been developed ; see, e.g., Abi Jaber et al. (2019); Tomas and Rosenbaum (2021). In Aichinger and Desmettre (2021), the authors analyze the Merton portfolio problems with power utility for a class of multivariate Volterra Heston models that features both inter-asset correlations and correlation between a stock and its volatility. In the present paper, we solve the Merton portfolio problem for investors with power and exponential utility functions, within the class of the so-called fake stationary multivariate affine Volterra models.
Main contributions. Building upon recent developments in volatility and Volterra models (Gnabeyeu et al., 2024, a, b) and motivated by recent works and advances on multivariate Volterra volatility modeling Abi Jaber et al. (2019); Tomas and Rosenbaum (2021); Aichinger and Desmettre (2021), the primary objective of this paper is to advance the literature on utility maximization along two main directions:
  • (i)
    We introduce a class of multivariate fake stationary affine Volterra stochastic volatility models that capture key stylized features of financial markets, including heterogeneous roughness across assets, possibly stochastic inter-asset correlations, and leverage effects namely, dependence between asset returns and their respective volatilities while maintaining a consistent modeling framework across time scales, from short to long maturities.
  • (ii)
    This model preserve analytical tractability, thereby enabling explicit characterization of the optimal investment strategy for the Merton’s problem, despite the intrinsic challenges posed by multivariate non-Markovian dynamics.
Organization of the Work. The outline of the paper is as follows: Section 2 gives an overview of the model which is needed throughout the paper: We introduce the multi asset financial market, where volatility is modeled by a multivariate class of fake stationary Volterra square root process, and we state the optimization problem. For such a market model we consider in section 3 two different approaches to solve the Merton portfolio problem. More precisely, in section 3 we investigate the classical problem of maximizing the expected utility of terminal wealth in a multi-asset fake stationary Volterra–Heston volatility market, for both power and exponential utility preferences. We first adapt the martingale distortion transformation used in Fouque and Hu (2019); Han and Wong (2020) to the (degenerate) multivariate case. However, as it is pointed out in Abi Jaber et al. (2021); Aichinger and Desmettre (2021), this only works if the correlation structure is highly degenerate. Inspired by the techniques used in Bäuerle and Li (2013); Aichinger and Desmettre (2021), we then provide a solution for the Merton portfolio optimization problem for a more general correlation structure using a verification argument. In Section 4, we demonstrate the practical implications of our findings through numerical experiments based on a two-dimensional fake stationary rough Heston model. Finally, Section 5 is devoted to the proofs of the main results.
Notations.
$\bullet$ Denote $\mathbb{T}=[0,T]\subset\mathbb{R}_{+}$, ${\rm Leb}_{d}$ the Lebesgue measure on $(\mathbb{R}^{d},{\cal B}or(\mathbb{R}^{d}))$, $\mathbb{H}:=\mathbb{R}^{d},$ etc.
$\bullet$ $\mathbb{X}:=C([0,T],\mathbb{H})(\text{resp.}\quad C_{0}([0,T],\mathbb{H}))$ denotes the set of continuous functions(resp. null at 0) from $[0,T]$ to $\mathbb{H}$ and ${\cal B}or(C_{d})$ denotes the Borel $\sigma$-field of ${C}_{d}$ induces by the $\sup$-norm topology.
$\bullet$ For $p\in(0,+\infty)$, $L_{\mathbb{H}}^{p}(\mathbb{P})$ or simply $L^{p}(\mathbb{P})$ denote the set of $\mathbb{H}$-valued random vectors $X$ defined on a probability space $(\Omega,{\cal A},\mathbb{P})$ such that $\|X\|_{p}:=(\mathbb{E}[\|X\|_{\mathbb{H}}^{p}])^{1/p}<+\infty$.
$\bullet$ Let $\mathcal{M}$ denote the space of all $(\mathbb{R}_{+},{\cal B}or(\mathbb{R}))$-measurable functions $m$ on $\mathbb{R}_{+}$ such that the restriction $\mu|_{[0,T]}$, for any $T>0$, is a $\mathbb{R}$-valued finite measure (i.e. the restriction $m|_{[0,T]}$ with $T>0$ is well-defined). For $m\in\mathcal{M}$ and a compact set $E\subset\mathbb{R}_{+}$, we define the total variation of $m$ on $E$ by:
$|m|(E):=\sup\left\{\sum_{j=1}^{N}|m(E_{j})|:\{E_{j}\}_{j=1}^{N}\text{ is a finite measurable partition of }E\right\}.$
We assume that the set of measure $m\in\mathcal{M}$ on $\mathbb{R}_{+}$ is of locally bounded variation.
$\bullet$ Convolution between a function and a measure. Let $f:(0,T]\to\mathbb{R}$ be a measurable function and $m\in\mathcal{M}$. Their convolution (whenever the integral is well-defined) is defined by
$(f*m)(t)=\int_{[0,t)}f(t-s)\,dm(s)=\int_{[0,t)}f(t-s)\,m(ds)=(f\stackrel{{\scriptstyle m}}{{*}}\mathbf{1})_{t},\quad t\in(0,T].$ (1.1)
$\bullet$ $X\perp\!\!\!\perp Y$ stands for independence of random variables, vectors or processes $X$ and $Y$.
$\bullet$ For a measurable function $\varphi:\mathbb{R}^{+}\to\mathbb{R}$, $\forall p\geq 1,$ we denote:
$\|\varphi\|^{p}_{L^{p}([0,T])}:=\int_{0}^{T}|\varphi(u)|^{p}\,du,\;\displaystyle\|\varphi\|_{\infty}=\|\varphi\|_{\sup}:=\sup_{u\in\mathbb{R}^{+}}|\varphi(u)|\;\text{and}\;\displaystyle\|\varphi\|_{\infty,T}=\|\varphi\|_{\sup,T}:=\sup_{u\in[0,T]}|\varphi(u)|.$
$\bullet$ $\Gamma(a)=\int_{0}^{+\infty}u^{a-1}e^{-u}\,du,\quad a>0,\quad\text{and}\quad B(a,b)=\int_{0}^{1}u^{a-1}(1-u)^{b-1}\,du,\quad a,b>0.$ We set $\mathbb{R}_{+}=[0,+\infty)$, $\mathbb{R}_{-}=(-\infty,0]$.
$\bullet$ Let $[0,T]$ be a finite time horizon, where $T<\infty$. Given a complete probability space $(\Omega,{\mathcal{F}},\mathbb{P})$ and a filtration $\mathbb{F}=({\mathcal{F}}_{t})_{t\geq 0}$ satisfying the usual conditions (We equip $(\Omega,{\mathcal{F}},\mathbb{P})$ with a right-continuous, $\mathbb{P}$-complete filtration $\mathbb{F}$), we denote by
$L^{\infty}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ $=\left\{Y:\Omega\times[0,T]\mapsto\mathbb{R}^{d},\;\mathbb{F}-\text{prog.\penalty 10000\ measurable and bounded a.s.}\right\}$
$L^{p}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ $=\left\{Y:\Omega\times[0,T]\mapsto\mathbb{R}^{d},\;\mathbb{F}-\text{prog.\penalty 10000\ measurable s.t.\penalty 10000\ }\mathbb{E}\Big[\int_{0}^{T}|Y_{s}|^{p}ds\Big]<\infty\right\}$
${\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ $=\left\{Y:\Omega\times[0,T]\mapsto\mathbb{R}^{d},\;\mathbb{F}-\text{prog.\penalty 10000\ measurable s.t.\penalty 10000\ }\sup_{t\leq T}|Y_{t}(w)|<\infty\mbox{ a.s.}\right\}$
${\mathbb{S}^{p}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ $=\left\{Y:\Omega\times[0,T]\mapsto\mathbb{R}^{d},\;\mathbb{F}-\text{prog.\penalty 10000\ measurable s.t.\penalty 10000\ }\mathbb{E}\Big[\sup_{0\leq t\leq T}|Y_{t}|^{p}\Big]<\infty\right\}.$
Here $|\cdot|$ denotes the Euclidian norm on $\mathbb{R}^{d}$. Classically, for $p\in(1,\infty)$, we define $L^{p,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ as the set of progressive processes $Y$ for which there exists a sequence of increasing stopping times $\tau_{n}\uparrow\infty$ such that the stopped processes $Y^{\tau_{n}}$ are in $L^{p}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ for every $n\geq 1$, and we recall that it consists of all progressive processes $Y$ s.t. $\int_{0}^{T}|Y_{t}|^{p}dt$ << $\infty$, a.s. Likewise for $\mathbb{S}^{p,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. To unclutter notation, we write $L^{p,loc}_{\mathbb{F}}([0,T])$ instead of $L^{p,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ when the context is clear.
$\bullet$ We will use the matrix norm $|A|=\operatorname{tr}(A^{\top}A)$ in this paper.
Our problem is defined under a given complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, with a filtration $\mathbb{F}=\{\mathcal{F}_{t}\}_{0\leq t\leq T}$ satisfying the usual conditions, supporting a $2d$-dimensional Brownian motion $(B,B^{\top})$ for $d\geq 1$. The filtration $\mathbb{F}$ is not necessarily the augmented filtration generated by $(B,B^{\top})$ ; thus, it can be a strictly larger filtration. Here $\mathbb{P}$ is a real-world probability measure from which a family of equivalent probability measures can be generated.
2 Preliminaries: Multivariate fake stationary affine Volterra models
Fix $T>0$, $d\in\mathbb{N}$. We let $K=\operatorname{diag}(K_{1},\ldots,K_{d})$ be diagonal with scalar kernels $K_{i}\in L^{2}([0,T],\mathbb{R})$ on the diagonal, $\varphi=\operatorname{diag}(\varphi^{1},\ldots,\varphi^{d})$, $\nu=\operatorname{diag}(\nu_{1},\ldots,\nu_{d})$, $\varsigma=\operatorname{diag}(\varsigma^{1},\ldots,\varsigma^{d})$ with $\varsigma^{i}$ a (locally) bounded Borel function and $D:=-\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})\in\mathbb{R}^{d\times d}$. Let $V=(V^{1},\ldots,V^{d})^{\top}$ be the following $\mathbb{R}^{d}_{+}$–valued scaled Volterra square–root process driven by an $d$-dimensional process $W=(W^{1},\ldots,W^{d})^{\top}$:
$V_{t}=\varphi(t)V_{0}+\int_{0}^{t}K(t-s)\big(\mu(s)+DV_{s}\big)ds+\int_{0}^{t}K(t-s)\nu\varsigma(s)\sqrt{\operatorname{diag}(V_{s})}dW_{s},\quad V_{0}\perp\!\!\!\perp W.$ (2.2)
Here $\mu:\mathbb{R}_{+}\to\mathbb{R}^{d}$, $W$ is a $d$-dimensional Wiener process. Note that the drift $b(t,x)=\mu(t)+Dx$ is clearly Lipschitz continuous in $x\in\mathbb{R}^{d}$, uniformly in $t\!\in\mathbb{T}_{+}$ and both the drift term $b$ and the diffusion coefficient $\sigma(t,x)=\nu\varsigma(t)\sqrt{\operatorname{diag}(x)}$ are of linear growth, i.e. there is a constant $C_{b,\sigma}>0$ such that
$\|b(t,x)\|+\||\sigma(t,x)|\|\leq C_{b,\sigma}(1+\|x\|),\quad\text{for all }t\in[0,T]\text{ and }x\in\mathbb{R}.$
We always work under the assumption below, which applies to the inhomogeneous Volterra equation (2.2).
Assumption 2.1 (On Volterra Equations with convolutive kernels).
Assume that $K$ is diagonal with scalar kernels $K_{i}$ on the diagonal for $i=1,\ldots,d$ that is completely monotone on $(0,\infty)$ and satisfy for any $T>0$,
  1. (i)
    Assume that the kernel $K_{i}$ is strictly positive and satisfy:
    • The integrability assumption: The following is satisfied for some $\widehat{\theta}_{i}\in(0,1]$.
      $(\widehat{\cal K}^{cont}_{\widehat{\theta}})\;\;\exists\,\widehat{\kappa_{i}}<+\infty,\;\forall\bar{\delta}\!\in(0,T],\;\widehat{\eta}(\delta):=\sup_{t\in[0,T]}\left[\int_{(t-\bar{\delta})^{+}}^{t}\thinspace K_{i}\big(t-u\big)^{2}du\right]^{\frac{1}{2}}\leq\widehat{\kappa}_{i}\,\bar{\delta}^{\,\widehat{\theta}_{i}}.$ (2.3)
    • The continuity assumption: $({\cal K}^{cont}_{\theta})\;\;\exists\,\kappa_{i}<+\infty,\;\exists\;\theta_{i}\in(0,1]\;\text{such that}\;\forall\,\bar{\delta}{\in(0,T)}$
      $({\cal K}^{cont}_{\theta})\;\forall\,\bar{\delta}{\,\in(0,T)},\;\eta(\bar{\delta}):=\sup_{t\in[0,T]}\left[\int_{0}^{t}|K_{i}(\big(s+\delta)\wedge T\big)-K_{i}(s)|^{2}ds\right]^{\frac{1}{2}}\leq\kappa_{i}\,\bar{\delta}^{\theta_{i}}.$ (2.4)
  2. (ii)
    Finally, assume that $V_{0}^{i}\in L^{p}(\mathbb{P})$ for some suitable $p\in(0,+\infty)$, such that the process $t\to v_{0}^{i}(t)=V_{0}\varphi^{i}(t)$ is absolutely continuous and $(\mathcal{F}_{t})$-adapted. Moreover, for some $\delta_{i}>0$, for any $p>0$,
    $\mathbb{E}\,\!\Big(\sup_{t\in[0,T]}|v_{0}^{i}(t)|^{p}\Big)<+\infty,\quad\mathbb{E}\!\big[\,|v_{0}^{i}(t^{\prime})-v_{0}^{i}(t)|^{p}\,\big]\leq C_{T,p}\Big(1+\mathbb{E}\,\big[\sup_{t\in[0,T]}|v_{0}^{i}(t)|^{p}\big]\Big)|t^{\prime}-t|^{\delta_{i}p}.$
Remark: For $i=1,\ldots,d$ , as $K_{i}$ is completely monotone on $(0,\infty)$ and not identically zero, we have that $K_{i}$ is nonnegative, not identically zero, non-increasing and continuous on $(0,\infty)$, and it follows from (B. Gripenberg and Saavalainen, 1990, Theorem 5.5.4) that $K_{i}$ has a resolvent of the first kind $r_{i}$ which is nonnegative and non-increasing in the sense that $s\mapsto r_{i}([s,s+t])$ is non-increasing for all $t\geq 0$.
In the case of $\alpha$- fractional kernel (corresponding to $K_{i}=K_{\alpha_{i}}$ with $\alpha_{i}\in[\frac{1}{2},1)$), by Gnabeyeu et al. (b) Equation (2.2) admits at least a unique-in-law positive weak solution as a scaling limit of a sequence of time-modulated Hawkes processes with heavy-tailed kernels in a nearly unstable regime. Moreover, under assumption 2.1 for some $p>0$, a solution $t\mapsto V_{t}$ to Equation (2.2) starting from $V_{0}$ has a $\big(\delta\wedge\vartheta\wedge\frac{\beta-1}{2\beta}-\eta\big)$-Hölder pathwise continuous modification on $\mathbb{R}_{+}^{d}$ for sufficiently small $\eta>0$ and satisfying (among other properties),
$\forall\,T>0,\;\exists\,C_{{}_{T,p}}>0,\quad\big\|\sup_{t\in[0,T]}\|V_{t}\|\big\|_{p}\leq C_{{}_{T,p}}\left(1+\big\|\sup_{t\in[0,T]}\|\varphi(t)V_{0}\|\big\|_{p}\right).$ (2.5)
Note that under our assumptions, if $p>0$ and $\mathbb{E}[\|\varphi(t)V_{0}\|^{p}]<+\infty$ for every $t\geq 0$, then by (2.5), $\mathbb{E}[\sup_{t\in[0,T]}\|V_{t}\|^{p}]<C_{T}(1+\mathbb{E}[\sup_{t\in[0,T]}\|\varphi(t)V_{0}\|^{p}])<+\infty$ for every $T>0$. Combined with the linear growth in Assumption 2.1(ii) $\||\sigma(t,x)|\|\leq C^{\prime}_{T}(1+\|x\|)$ for $t\in[0,T]$, this implies $\mathbb{E}[\sup_{t\in[0,T]}\||\sigma(t,X_{t})|\|^{p}]<C^{\prime}_{T}(1+\mathbb{E}[\sup_{t\in[0,T]}\|\varphi(t)V_{0}\|^{p}])<+\infty$ for every $T>0$, enabling the unrestricted use of both regular and stochastic Fubini’s theorems. Sufficient conditions for interchanging the order of ordinary integration (with respect to a finite measure) and stochastic integration (with respect to a square integrable martingale) are provided in (Kailath et al., 1978, Thm.1), and further details can be found in (Protter, 2005, Thm. IV.65), (Walsh, 1986, Theorem 2.6), (Veraar, 2012, Theorem 2.6).
Remark 2.1.
This covers, for instance, constant non-negative kernels, fractional kernels of the form $\frac{t^{\alpha-1}}{\Gamma(\alpha)}\mathbf{1}_{\mathbb{R}_{+}}$ with $\alpha\in(\frac{1}{2},1]$, exponentially decaying kernels ${\rm e}^{-\beta t}$ with $\beta>0$ and more generally the gamma kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}e^{-\beta t}\mathbf{1}_{\mathbb{R}_{+}}$ with $\alpha\in\left(\tfrac{1}{2},1\right]$ and $\beta\geq 0$ ( see e.g. (Gnabeyeu and Pagès, 2025, Propositions 6.1 and 6.3)).
The roughness of the volatility paths is determined by the parameter $\alpha$ linked to the Hurst parameter $H$ via the relation $\alpha=H+\frac{1}{2}$. For $\alpha\rightarrow 1$ we recover the classical markovian square root process.
2.1 Stabilizer and fake stationarity regimes.
Definition 2.2 (Fake Stationarity Regimes).
Let $(V_{t})_{t\geq 0}$ be a solution to the scaled Volterra equation (2.2) starting from any $V_{0}\in L^{2}(\mathbb{P})$. Then, the process $(V_{t})_{t\geq 0}$ exhibit a fake stationary regime of type I in the sense of Pagès (2024); Gnabeyeu and Pagès (2025) if it has constant mean and variance over time i.e.:
$\forall\,t\geq 0,\quad\mathbb{E}[V_{t}]=\textit{c}^{\text{ste}}\quad\mbox{and}\quad\text{Var}(V_{t})=\textit{c}^{\text{ste}}=v_{0}\in\mathbb{R}_{+}^{d}.$ (2.6)
For every $\lambda\!\in\mathbb{R}$, the resolvent or Solvent core $R_{\lambda}$ associated to a real-valued kernel $K$, known as the $\lambda$-resolvent of $K$ is defined as the unique solution – if it exists – to the deterministic Volterra equation
$\forall\,t\geq 0,\quad R_{\lambda}(t)+\lambda\int_{0}^{t}K(t-s)R_{\lambda}(s)ds=1.$ (2.7)
or, equivalently, written in terms of convolution, $R_{\lambda}+\lambda K*R_{\lambda}=1$ and admits the formal Neumann series expansion $R_{\lambda}=\mbox{\bf 1}*\big(\sum_{k\geq 0}(-1)^{k}\lambda^{k}K^{k*}\big)$ where $K^{k*}$ denotes the $k$-th convolution of $K$ with the convention, $K^{0*}=\delta_{0}$ (Dirac mass at 0).
Remark If $K$ is regular enough (say continuous) the resolvent $R_{\lambda}$ is differentiable and one checks that $f_{\lambda}=-R^{\prime}_{\lambda}$ satisfies for every $t>0$, $f_{\lambda}(t)+\lambda\big(R_{\lambda}(0)K(t)-K*f_{\lambda}(t)\big)=0$ that is $f_{\lambda}$ is solution to the equation
$f_{\lambda}+\lambda K*f_{\lambda}=\lambda K\quad\text{and reads}\quad f_{\lambda}=\sum_{k\geq 1}(-1)^{k}\lambda^{k}K^{k*},\quad K^{0*}=\delta_{0}.$ (2.8)
Example 2.3.
Denote by $E_{\alpha}$ the standard Mittag-Leffler function. For the $\alpha$-fractional kernels defined in Remark 2.1 the identity $K_{\alpha}*K_{\alpha^{\prime}}=K_{\alpha+\alpha^{\prime}}$ holds for $t\geq 0$ so that
$R_{\alpha,\lambda}(t)=\sum_{k\geq 0}(-1)^{k}\frac{\lambda^{k}t^{\alpha k}}{\Gamma(\alpha k+1)}=E_{\alpha}(-\lambda t^{\alpha}),\;\text{and }\;f_{\alpha,\lambda}(t)=-R^{\prime}_{\alpha,\lambda}(t)=\lambda t^{\alpha-1}\sum_{k\geq 0}(-1)^{k}\lambda^{k}\frac{t^{\alpha k}}{\Gamma(\alpha(k+1))}.$
We will always work under the following assumption.
Assumption 2.2 ($\lambda$-resolvent $R_{\lambda}$ of the kernel).
For $i=1,\cdots,d$, we assume that the $\lambda_{i}$-resolvent $R_{\lambda_{i}}$ of the kernel $K_{i}$ satisfies the following for every $\lambda_{i}>0$:
[…]
Remark: Under the assumption $({\cal K})$, $f_{\lambda_{i}}$ is a $(1-a_{i})$-sum measure, i.e., $\int_{0}^{+\infty}f_{\lambda_{i}}(s)\,ds=1-a_{i}$. Furthermore, $\lim_{t\to+\infty}\int_{0}^{t}f_{\lambda_{i}}(t-s)\mu^{i}(s)ds=\mu_{\infty}^{i}$ and $\lim_{t\to+\infty}\varphi^{i}(t)-(f_{\lambda_{i}}*\varphi^{i})(t)=\varphi_{\infty}^{i}\,a_{i}.$ (see (Gnabeyeu and Pagès, 2025, Lemma 3.1)). Finally, if $f_{\lambda_{i}}=-R^{\prime}_{\lambda_{i}}>0\text{ for }t>0$, then $f_{\lambda_{i}}$ is a probability density in which case, $R_{\lambda_{i}}$ is non-increasing. This is in particular the case for the Mittag-Leffler density function $f_{\alpha_{i},\lambda_{i}}$ for $\alpha_{i}\in(\frac{1}{2},1)$, in which case $f_{\alpha_{i},\lambda_{i}}$ is a completely monotonic function (hence convex), decreasing to 0 while $1-R_{\alpha_{i},\lambda_{i}}$ is a Bernstein function (see e.g. (Gnabeyeu and Pagès, 2025, Proposition 6.1)). The Proposition below shows what are the consequences of the three constraints in equation (2.6).
Proposition 2.4 (Fake stationary Volterra square root process.).
Let $(V_{t})_{t\geq 0}$ be a solution to the scaled Volterra square root equation in its form (2.2) starting from any random variable $V_{0}\in L^{2}(\Omega,\mathcal{F},\mathbb{P})$. Then, a necessary and sufficient condition for the relations (2.6) to be satisfied is that for $i=1,\ldots,d$
$\ \mathbb{E}[V_{0}^{i}]=\frac{1-a_{i}}{1-a_{i}\varphi_{\infty}^{i}}\frac{\mu_{\infty}^{i}}{\lambda_{i}}:=x_{\infty}^{i}\qquad\text{and}\qquad\forall\,t\geq 0,\qquad\varphi^{i}(t)=1-\lambda_{i}\int_{0}^{t}K_{i}(t-s)\left(\frac{\mu^{i}(s)}{\lambda_{i}x_{\infty}^{i}}-1\right)\,\,\mathrm{d}s.$ (2.10)
[…]
and the couple $(v_{0}^{i},\varsigma^{i}(t))$, where $v_{0}^{i}=\text{Var}(V_{0}^{i})$ must satisfy the functional equation:
$\textit{($E_{\lambda_{i},c_{i}}$)}:\;\forall\,t\geq 0,\;c_{i}\lambda_{i}^{2}\big(1-(\varphi^{i}(t)-(f_{\lambda_{i}}*\varphi^{i})_{t})^{2}\big)=(f_{\lambda_{i}}^{2}*\varsigma^{i2})(t)\;\textit{where}\;c_{i}=\frac{v_{0}^{i}}{\nu_{i}^{2}x_{\infty}^{i}}\;\textit{i.e.}\;\varsigma^{i}=\varsigma^{i}_{\lambda_{i},c_{i}}.$ (2.12)
Proof : This is a straightforward extension to the multi-dimensional setting of (Gnabeyeu and Pagès, 2025, Proposition 3.4 and Theorem 3.5) (see also (Gnabeyeu et al., a, Proposition 4.2 and 4.4)).
Definition 2.5.
We will call the stabilizer (or corrector) of the scaled stochastic Volterra equation (2.2) the (locally) bounded Borel function $\varsigma=\operatorname{diag}(\varsigma^{1},\ldots,\varsigma^{d})$ where $\varsigma^{i}$ is a solution(if any) to the functional equation ($E_{\lambda_{i},c_{i}}$) in (2.12) for $i=1,\ldots,d$.
Example 2.6.
Within the setting $\varphi^{i}(t)=\varphi^{i}(0)=1$ for all $t\geq 0$ and $K_{i}$ the $\alpha$-fractional kernel defined in Remark 2.1 and Example 2.3 with $\alpha_{i}\in\left(\frac{1}{2},1\right)$, we have $\lim_{t\to+\infty}R_{\alpha_{i},\lambda_{i}}=0$. Setting $a_{k}=\frac{1}{\Gamma(\alpha k+1)},b_{k}=\frac{1}{\Gamma(\alpha(k+1))},\;k\geq 0$, then the stabilizer $\varsigma=\varsigma_{\alpha_{i},\lambda_{i},c_{i}}$ exists as a non-negative, non-increasing concave function, on $(0,+\infty)$ (see (Pagès, 2024, Sections 5.1 and 5.2 ), (Gnabeyeu and Pagès, 2025, Sections 5.1 and 5.2 )), such that:
$\varsigma^{2}_{\alpha_{i},\lambda_{i},c_{i}}(t)=c_{i}\lambda_{i}^{2-\frac{1}{\alpha_{i}}}\varsigma_{\alpha_{i}}^{2}(\lambda_{i}^{\frac{1}{\alpha_{i}}}t)$ where $\varsigma_{\alpha_{i}}^{2}(t):=2\,t^{1-\alpha_{i}}\sum_{k\geq 0}(-1)^{k}c_{k}t^{\alpha_{i}k}$ and the coefficients $(c_{k})_{k\geq 0}$ are defined by the recurrence formula $c_{0}=\frac{\Gamma(\alpha)^{2}}{\Gamma(2\alpha-1)\Gamma(2-\alpha)}$ and for every $k\geq 1$
$c_{k}=\frac{\Gamma(\alpha)^{2}\Gamma(\alpha(k+1))}{\Gamma(2\alpha-1)\Gamma(\alpha k+2-\alpha)}\left[(a*b)_{k}-\alpha(k+1)\sum_{\ell=1}^{k}B\big(\alpha(\ell+2)-1,\alpha(k-\ell-1)+2\big)(b^{*2})_{\ell}c_{k-\ell}\right].$ (2.13)
where for two sequences of real numbers $(u_{k})_{k\geq 0}$ and $(v_{k})_{k\geq 0}$, the Cauchy product is defined as $(u*v)_{k}=\sum_{\ell=0}^{k}u_{\ell}v_{k-\ell}$ and $B(a,b)=\int_{0}^{1}u^{a-1}(1-u)^{b-1}du$ denoting the beta function.
Moreover, $\left(\liminf_{k}\left(|c_{k}|^{1/k}\right)\right)^{-1/\alpha}=\infty$, $\varsigma_{\alpha_{i},\lambda_{i},c_{i}}(0)=0$ and $\lim_{t\to+\infty}\varsigma_{\alpha_{i},\lambda_{i},c_{i}}(t)=\frac{\sqrt{c_{i}}\lambda_{i}}{\|f_{\alpha_{i},\lambda_{i}}\|_{L^{2}(\text{Leb}_{1})}}$.
Set $E_{D,c}=\bigcup_{i=1}^{d}E_{\lambda_{i},c_{i}}$. From now on, we will assume that there exists a unique positive bounded Borel solution $\varsigma=\varsigma_{D,c}$ on $(0,+\infty)$ of the system of equation $(E_{D,c})$ so that, the corresponding time-inhomogeneous Volterra square root equation (2.2) is refered to as a Multivariate Stabilized Volterra Cox-Ingersoll-Ross (CIR) equation or as a Multivariate fake stationary Volterra CIR equation if, in addition, equation (2.10) holds.
2.2 Formulation of the stochastic Market model
We consider a financial market on $[0,T]$ on some filtered probability space $(\Omega,{\mathcal{F}},\mathbb{F}:=({\mathcal{F}}_{t})_{t\geq 0},\mathbb{P})$ with $d+1$ securities, consisting of a bond and $d$ stocks. The non–risky asset $S^{0}$ satisfies the (stochastic) ordinary differential equation:
$dS^{0}_{t}=S^{0}_{t}r(t)dt,$
with a time-dependent deterministic short risk-free rate $r:\mathbb{R}_{+}\to\mathbb{R}$, and $d$ risky assets (stock or index) whose return vector process $(S_{t})_{t\geq 0}=(S_{t}^{1},\ldots,S_{t}^{d})_{t\geq 0}$ is defined via the dynamics given by the vector-stochastic differential equation (SDE):
$dS_{t}=\operatorname{diag}(S_{t})\big[\big(r(t){\mathbb{1}_{d}}+\sigma_{t}\lambda_{t}\big)dt+\sigma_{t}dB_{t}\big],$ (2.14)
driven by a $d$-dimensional Brownian motion $B$, with a $d\times d$-matrix valued continuous stochastic volatility process $\sigma$ whose dynamics is driven by (2.2) and a $\mathbb{R}^{d}$-valued continuous stochastic process $\lambda$, called market price of risk. Here $\mathbb{1}_{d}$} denotes the vector in $\mathbb{R}^{d}$ with all components equal to $1$ and the correlation structure of $W$ with $B$ is given by
$W^{i}=\rho_{i}B^{i}+\sqrt{1-\rho_{i}^{2}}B^{\perp,i}=\Sigma_{i}^{\top}B_{t}+\sqrt{1-\Sigma_{i}^{\top}\Sigma_{i}}B^{\perp,i}_{t},\quad i=1,\ldots,d,$ (2.15)
for some $(\rho_{1},\ldots,\rho_{d})\in[-1,1]^{d}$, where $(0,\ldots,\rho_{i},\ldots,0)^{\top}:=\Sigma_{i}\in\mathbb{R}^{d}$ is such that $\Sigma_{i}^{\top}\Sigma_{i}\leq 1$, and $B^{\perp}$ == $(B^{\perp,1},\ldots,B^{\perp,d})^{\top}$ is an dd–dimensional Brownian motion independent of $B$. The correlation $\rho_{i}$ between stock price $S^{i}$ and variance $V^{i}$ is assumed constant. Note that $d\langle W^{i}\rangle_{t}=dt$ but $W^{i}$ and $W^{j}$ can be correlated, hence $W$ is not necessarily a Brownian motion.
Observe that processes $\lambda$ and $\sigma$ are $\mathbb{F}$-adapted, possibly unbounded, but not necessarily adapted to the filtration generated by $W$. We point out that $\mathbb{F}$ may be strictly larger than the augmented filtration generated by $B$ and $B^{\perp}$ as we deal with weak solutions to stochastic Volterra equations.
We assume that $\sigma$ in (2.14) is given by $\sigma=\sqrt{\operatorname{diag}(V)}$, where the $\mathbb{R}^{d}_{+}$–valued scaled process $V$ is defined in (2.2) with $\varsigma=\varsigma_{D,c}$ and Equation (2.10) holds true. We will be chiefly interested in the case where $\lambda_{t}$ is linear in $\sigma_{t}$. More specifically, the the market price of risk (risk premium) is assumed to be in the form $\lambda$ == $\big(\theta_{1}\sqrt{V^{1}},\ldots,\theta_{d}\sqrt{V^{d}}\big)^{\top}$, for some constant $\theta_{i}\geq 0$, so that the dynamics for the stock prices (2.14) reads following Kraft (2005); Abi Jaber et al. (2019)
$dS^{i}_{t}=S^{i}_{t}\left(r(t)+\theta_{i}V^{i}_{t}\right)dt+S^{i}_{t}\sqrt{V^{i}_{t}}dB^{i}_{t},\quad i=1,\ldots,d.$ (2.16)
Since $S$ is fully determined by $V$, the existence of $S$ readily follows from that of $V$. In particular, weak existence of Hölder pathwise continuous solution $V$ of (2.2) such that (2.5) holds is established under suitable assumptions on the kernel $K$ and specifications $g_{0}$ as shown in the following remark.
We state the following existence and uniqueness result from Gnabeyeu et al. (b) which is extended to the multi-dimensional setting.
Theorem 2.7.
((Gnabeyeu et al., b, Theorem 3.1 and Remark on Theorem 3.2)). Under Assumption 2.1, the stochastic Volterra equation (2.16)-(2.2) has a unique in law continuous $\mathbb{R}^{d}_{+}\times\mathbb{R}^{d}_{+}$-valued weak solution $(S,V)$ for any initial condition $(S_{0},V_{0})\in\mathbb{R}^{d}_{+}\times\mathbb{R}^{d}_{+}$ defined on some filtered probability space $(\Omega,\mathcal{F},(\mathcal{F})_{t\geq 0},\mathbb{P})$ such that
$\sup_{t\leq T}\mathbb{E}\left[\|V_{t}\|^{p}\right]<\infty,\quad p>0.$ (2.17)
From now on, we set $g_{0}(t):=\varphi(t)V_{0}+\int_{0}^{t}K(t-s)\mu(s)ds$ and $dZ_{t}=DV_{t}dt+\nu\sqrt{\operatorname{diag}(V_{t})}dW_{t},\;\forall t\geq 0$ so that Equation (2.2) reads
$V_{t}=g_{0}(t)+\int_{0}^{t}K(t-s)DV_{s}ds+\int_{0}^{t}K(t-s)\nu\varsigma(s)\sqrt{\operatorname{diag}(V_{s})}dW_{s}=g_{0}(t)+\int_{0}^{t}K(t-s)dZ_{s}.$ (2.18)
Finally, we consider the $\mathbb{R}^{d}$-valued process for $s\geq t,$
$g_{t}(s)=g_{0}(s)+\int_{0}^{t}K(s-u)\big(DV_{u}du+\nu\varsigma(s)\sqrt{\operatorname{diag}(V_{u})}dW_{u}\big)=g_{0}(s)+\int_{0}^{t}K(s-u)dZ_{u}.$ (2.19)
One notes that for each, $s\leq T$, $(g_{t}(s))_{t\leq s}$ is the adjusted forward process
$g_{t}(s)$ $=\;\mathbb{E}\Big[V_{s}-\int_{t}^{s}K(s-u)DV_{u}du{\ \Big|\ }{\mathcal{F}}_{t}\Big].$ (2.20)
This adjusted forward process is commonly used (see, e.g., Abi Jaber et al. (2021)) to elucidate the affine structure of affine Volterra processes with continuous trajectories.
The process in (2.18) is non-Markovian and non-semimartingale in general. Note that our model (2.16)-(2.15)-(2.2) features correlation between the stocks and between a stock and its volatility. Moreover, the methodology developed in this paper, and hence the results obtained, remain valid if the matrix $D$ in (2.2) is not assumed to be diagonal, but only satisfies
$D\in\mathbb{R}^{d\times d},\qquad D_{ij}\geq 0\ \text{for }i\neq j.$
This also provides an extension to the inhomogeneous setting of the models considered in Abi Jaber et al. (2019, 2021); Aichinger and Desmettre (2021).
3 Merton’s portfolio problem: Utility maximisation
[…]
We consider the classical problem of maximizing the expected utility of terminal wealth. A portfolio strategy $\alpha_{t}=(\alpha_{t,1},\ldots,\alpha_{t,d})^{\top}$ is an $\mathbb{R}^{d}$-valued, $\mathbb{F}$-progressively measurable process, where $\alpha_{t,k}$ represents the proportion of wealth invested in asset $k$ at time $t$. Under a fixed portfolio strategy, the corresponding wealth process $(X_{t}^{\alpha})_{t\geq 0}$ depend on $V$ in (2.2) and has a certain dynamic to be specified latter. By $\mathcal{A}$ we denote the set of admissible portfolio or investment strategies i.e. the set of all $\mathbb{F}$-progressively measurable processes $(\alpha_{t})_{t\in[0,T]}$ valued in the Polish space $\mathbb{R}^{d}$.
To ease notation burden, whenever the context is clear, we simply write $X$, instead of $X^{\alpha}$, as the wealth process under $\alpha\in{\mathcal{A}}$ with initial condition $X_{0}=x_{0}>0$ and $V_{0}>0$.
Definition 3.1.
  1. 1.
    A utility function is a strictly increasing and strictly concave function $U:(0,\infty)\to\mathbb{R}\cup\{-\infty\},$ which is continuously differentiable on $(0,\infty)$.
[…]
  2. 2.
    An admissible strategy $\alpha\in\mathcal{A}$ is said to be optimal (for terminal wealth) if it maximizes $\alpha\longmapsto\mathbb{E}\big[U(X_{T}^{\alpha})\big]$ over all admissible strategies in $\mathcal{A}$. That is the portfolio strategy $\alpha^{*}$ for which the supremum is attained
The investor’s goal in the Merton problem is to find an optimal strategy so as to maximize the expected utility of terminal wealth. Specifically, given the utility function $U$ on $(0,\infty)$ and starting from an initial capital $x_{0}>0$, the objective of the agent is
$\mathcal{V}(x_{0},V_{0}):=\sup_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}\!\left[U\!\left(X_{T}^{\alpha}\right)\right],\;\text{given}\;x_{0}\;\text{and}\;V_{0}.$ (3.21)
with $X^{\alpha}$ the wealth controlled by $\alpha\in\mathcal{A}$, starting from $x_{0}$ at time 0 and $\mathcal{A}$ the subset of controls $\alpha\in\mathcal{A}$ such that the family $\left\{U\!\left(X_{\tau}^{\alpha}\right):\tau\in[0,T]\right\}$ is uniformly integrable. The preferences of the agent are thus described by the utility function $U$. Mathematically, we aim at identifying the optimal value associated with the utility maximization problem defined over a finite horizon $T$ by
$\mathcal{V}_{t}(x,v):=\operatorname*{ess\,sup}_{\alpha\in\mathcal{A}_{t}}\mathbb{E}_{x,v}\!\left[U\!\left(X_{T}^{t,x,v,\alpha}\right)\mid\mathcal{F}_{t}\right],\qquad x>0,v>0,\;t\in[0,T],$ (3.22)
and the optimal strategy $\alpha^{\ast}$, given the preference described by the utility function $U(\cdot)$. This is a standard stochastic control problem. The set $\mathcal{A}_{t}$ is the class of all admissible strategies from t with zero being an absorbing state for $X_{t}^{\alpha}$ (bankruptcy). Here $\mathbb{E}_{x,v}$ denotes expectation under the conditional distribution given $X_{t}=x$ and $V_{t}=v$, and where the supremum is taken over all admissible portfolio strategies.
The $\mathbb{F}$-adapted process $\{\mathcal{V}_{t},\,0\leq t\leq T\}$ is a supermartingale. Moreover, there exists an optimal control $\alpha^{*}\in\mathcal{A}$ for (3.21) if and only if the martingale property holds, that is, the process $\{\mathcal{V}^{*}_{t},\,0\leq t\leq T\}$ for $\alpha^{*}$ is a martingale. The idea is that an admissible control (or startegy) is optimal333optimal strategy: No expected gain from deviation, this implies martingality if the associated value process is a martingale and for any other admissible control, it is a supermartingale. That is the classic martingale optimality principle, see, e.g., Hu et al. (2005), Pham (2009, Section 6.6.1) or Jeanblanc et al. (2012).
Definition 3.2 (Martingale optimality principle).
The Problem (3.21) can be solved by constructing a family of processes $\{J^{\alpha}_{t}\}_{t\in[0,T]}$, $\alpha\in{\mathcal{A}}$, satisfying the conditions:
  1. 1.
    $J^{\alpha}_{T}=U(X_{T})$ for all $\alpha\in{\mathcal{A}}$;
  2. 2.
    $J^{\alpha}_{0}$ is a constant, independent of $\alpha\in{\mathcal{A}}$;
  3. 3.
    $J^{\alpha}_{t}$ is a supermartingale for all $\alpha\in{\mathcal{A}}$, and there exists $\alpha^{*}\in{\mathcal{A}}$ such that $J^{\alpha^{*}}$ is a martingale.
A family of processes with the above properties can now be used to compare the expected utilities of an arbitrary strategy $\alpha\in{\mathcal{A}}$ and the strategy $\alpha^{*}$:
$\mathbb{E}[U(X_{T}^{\alpha})]=\mathbb{E}[J^{\alpha}_{T}]\leq J^{\alpha}_{0}=J^{\alpha^{*}}_{0}=\mathbb{E}[J^{\alpha^{*}}_{T}]=\mathbb{E}[U(X^{\alpha^{*}}_{T})]=\mathcal{V}(x_{0},V_{0}).$
where $X^{\alpha^{*}}$ is the wealth process under $\alpha^{*}$. Thus the strategy $\alpha^{*}$ is indeed our desired optimal portfolio strategy.
As stated before, Problem (3.21) seen as an optimization problem with state process $X^{\alpha}$ is non-Markovian and the standard stochastic control approach cannot be applied. Heuristically speaking, the non-Markovian and non-semimartingale characteristics of the multivariate fake stationary Volterra Heston model (2.16)-(2.2) are overcome in the degenerate correlation case by applying the martingale optimality principle and constructing an ansatz, which is inspired by the martingale distortion transformation and the exponential-affine representation (see e.g. (Gnabeyeu, 2026, Theorem A.4.)) in term of an auxiliary process denoted $\Gamma\equiv\Gamma(V_{t})$ (more precisely $\Gamma(g_{t})$). In the general correlation case, we provide a verification argument avoiding restrictions on the correlation structure linked to the martingale distortion transformation.
In fact, Fouque and Hu (2019) showed that if the Sharpe-ratio $\lambda$ in 2.14 is bounded and has bounded derivative, then the value process $\mathcal{V}_{t}$ can be expressed as $\mathcal{V}_{t}(x,v)=J_{t}(X_{t}^{\alpha}=x,V_{t}=v)$, where
$J_{t}(X_{t}^{\alpha},V_{t}):=F(t,X_{t}^{\alpha})\Gamma(V_{t})$ (3.23)
for the power utility case ($F(t,X_{t}^{\alpha})=U(X_{t}^{\alpha})=\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}$) even if the volatility process $V_{t}$ is non-Markovian. This approach is called the martingale distortion transformation and was first introduced in the seminal paper Zariphopoulou (2001) and later transferred to a non-Markovian setting with Hölder-type inequalities in Tehranchi (2004) for both power utility and exponential utility case (in which $F(t,X_{t}^{\alpha})=-\frac{1}{\gamma}\exp\Big(-\gamma e^{\int^{T}_{t}r(s)ds}X_{t}^{\alpha}\Big)$). The extension to the multi-asset case is straight forward in the case of a bounded risk premium and degenerate correlation structure(cf. Fouque and Hu (2019), Remark 2.5.).
We will provide concrete specifications of the process $\Gamma$ depending on the problem at hand. The semi-closed form exponential–affine representation (Gnabeyeu, 2026, Theorem A.4.) of the functional $\Gamma$ can be expressed in terms of the continuous solution $\psi$ of an associated time-inhomogeneous Riccati-type Volterra equation, which can be solved by well-known efficient numerical methods.
Note that, in contrast to Han and Wong (2020); Aichinger and Desmettre (2021), this representation is expressed in terms of the adjusted forward process, as defined in (2.19)–(2.20), rather than the forward process used in these works. We offer explicit solutions to the optimal portfolio policies that depend on the function $\psi$, solution of the above-mentionned time-dependent multivariate Riccati-Volterra equation. Let $\Lambda$ be defined as
$\Lambda_{t}^{i}=\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\sqrt{V^{i}_{t}},\quad i=1,\ldots,d,\quad 0\leq t\leq T,$ (3.24)
We will work under the following assumption,
Assumption 3.1.
Assume that there exists a solution $\psi\in C([0,T],\mathbb{R}^{d})$ to the above-mentioned inhomogeneous Riccati–Volterra equation satisfying the below appropriate boundedness condition i.e. such that
$\max_{1\leq i\leq d}\sup_{t\in[0,T]}\left(\theta_{i}^{2}+\nu_{i}^{2}\varsigma^{i}(t)^{2}\psi^{i}(T-t)^{2}\right)\leq\frac{a}{a(p)},$ (3.25)
holds for a sufficient large $p>1$, where the constant $a(p)$ is given by
$a(p)=\max\Big[p\left(2+|\Sigma|\right),{2(8p^{2}{-2p})\left(1+|{\Sigma}|^{2}\right)},{p\left(1+|{\Sigma}|^{2}\right)}\Big].$ (3.26)
and the constant $a>0$ is such that $\mathbb{E}\left[\exp\big(a\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\big)\right]<\infty$.
Remark on Assumption 3.1: Note that if Assumption 3.1 hold, then
$\mathbb{E}\Big[\exp\Big(a(p)\int_{0}^{T}\big(|\lambda_{s}|^{2}+\left|\Lambda_{s}\right|^{2}\big)ds\Big)\Big]\;<\;\infty,$ (3.27)
holds for some $p>1$ and a constant $a(p)$ given by (3.26).
In fact, under Assumption 3.1, we will have
$a(p)\left(|\lambda_{s}|^{2}+\left|\Lambda_{s}\right|^{2}\right)\;=\;a(p)\sum_{i=1}^{d}V_{s}^{i}\left(\theta_{i}^{2}+\nu_{i}^{2}\varsigma^{i}(s)^{2}\psi^{i}(T-s)^{2}\right)\;\leq\;a\sum_{i=1}^{d}V_{s}^{i},$ (3.28)
which implies that $\mathbb{E}\left[\exp\left(a(p)\int_{0}^{T}\left(|\lambda_{s}|^{2}+\left|\Lambda_{s}\right|^{2}\right)ds\right)\right]<\infty$.
Remark: Condition (3.25) concerns the risk premium constants $(\theta_{1},\ldots,\theta_{d})$. For a large enough constant $a>0$, from (Gnabeyeu, 2026, Theorem A.4.) ( with $\mathcal{M}\ni m(\,\mathrm{d}s):=a{\mathbb{1}_{d}}\,\delta_{0}(\,\mathrm{d}s)$), a sufficient condition ensuring $\mathbb{E}\big[\exp\big(a\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\big)\big]<\infty$ is the existence of a continuous solution $\tilde{\psi}$ on $[0,T]$ to the inhomogeneous Riccati–Volterra equation
$\tilde{\psi}^{i}(t)$ $=\;\int_{0}^{t}K_{i}(t-s)\Big(a+\big(D^{\top}\tilde{\psi}(s)\big)_{i}+\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(T-s)\tilde{\psi}^{i}(s))^{2}\Big)ds,\;\;i=1,\ldots,d,\;\;0\leq t\leq T.$ (3.29)
Power and exponential utility functions are widely adopted in the literature and display distinct risk-aversion properties. Specifically, the power utility function exhibits constant relative risk aversion (CRRA), whereas the exponential utility function is characterized by constant absolute risk aversion (CARA). Optimal strategies corresponding to these utility functions are presented in the sequel, within the framework of the model introduced in Section 2.
3.1 Optimal strategy for the power utility maximization problem
Here, we denote by $\pi_{t}$ the proportion of wealth invested in the risky assets $S$. An agent invests at any time $t$ the proportion $\pi_{t}$ of his wealth $X^{\pi}_{t}$ in the stocks $S$ of price’s vector $S_{t}$, and the remaining proportion $1-\pi_{t}^{\top}{\mathbb{1}_{d}}$ in a bond of price $S_{t}^{0}$ with interest rate $r(t)$. The notation $X^{\pi}_{t}$ emphasizes the dependence of the wealth on the strategy $\pi=(\pi_{t})_{t\geq 0}$. The portfolio strategy $\pi_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{d})$ is an $(\mathbb{R}^{d})^{*}$ valued, progressively measurable process, where $\pi_{t}^{k}$ represents the proportion of wealth invested into stock $k$ at time $t$. Assuming the model (2.14)– (2.16) for $S_{t}$, the dynamics of the controlled wealth process is given by
$\mathrm{d}X^{\pi}_{t}$ $=X^{\pi}_{t}\bigl(\pi_{t}^{\top}\big(r(t){\mathbb{1}_{d}}+\sigma(V_{t})\lambda_{t}\big)+(1-\pi_{t}^{\top}{\mathbb{1}_{d}})r(t)\bigr)\,\mathrm{d}t+X^{\pi}_{t}\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}$
$=X^{\pi}_{t}\bigl(r(t)+\pi_{t}^{\top}\sigma(V_{t})\lambda_{t}\bigr)\,\mathrm{d}t+X^{\pi}_{t}\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}=X_{t}^{\pi}(r(t)+\pi_{t}^{\top}{\operatorname{diag}(V_{t})}{\theta})dt+X_{t}\pi_{t}^{\top}\sqrt{\operatorname{diag}(V_{t})}dB_{t},$
where ${\theta}=(\theta_{1},\dots,\theta_{d})^{\top}$. The investor faces the portfolio constraint that, at any time $t$, $\pi_{t}$ takes values in a closed convex subset $\mathcal{A}\subset\mathbb{R}^{d}$ denoting the set of admissible portfolio strategies.
From now, we let $\alpha_{t}:=\sigma(V_{t})^{\top}\pi_{t}$ be the investment strategy. Then, under a fixed portfolio strategy, the wealth process $X^{\alpha}_{t}$, controlled by $\alpha$ is governed by:
$dX^{\alpha}_{t}=\big(r(t)+\alpha_{t}^{\top}\lambda_{t}\big)X^{\alpha}_{t}dt+\alpha_{t}^{\top}X^{\alpha}_{t}dB_{t},\;X_{0}=x_{0}>0.$ (3.30)
By solving the linear SDE (3.30), the wealth process admits the explicit representation
$X_{t}^{\alpha}=x_{0}\exp\!\left(\int_{0}^{t}\Big(r(s)+\alpha_{s}^{\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}\right|^{2}\Big)\,ds+\int_{0}^{t}\alpha_{s}^{\top}\,dB_{s}\right),\quad x_{0}\geq 0.$ (3.31)
By a solution to (3.30), we mean an $\mathbb{F}$-adapted process $X^{\alpha}$ satisfying (3.30) on $[0,T]$ with $\mathbb{P}$-a.s. continuous sample paths and such that
$\mathbb{E}\big[\sup_{t\leq T}|X^{\alpha}_{t}|^{p}\big]$ $<\;\infty\quad\text{for some }\quad p>1.$ (3.32)
The conditions under which we consider a strategy to be admissible is specified in the below definition.
Definition 3.3.
In the setting described above, we say that an investment strategy $\alpha(\cdot)$ is admissible if
  1. $(a)$
    The SDE (3.30) for the wealth process $(X_{t}^{\alpha})$ has a unique solution in terms of $(S,V,B)$; with $\mathbb{P}$-a.s. continuous paths and $X_{t}\geq 0$, $\forall\;t\in[0,T]$, $\mathbb{P}$-a.s.;
  2. $(b)$
    $\mathbb{E}[\frac{1}{\gamma}(X_{T}^{\alpha})^{\gamma}]<\infty$ for all $0<\gamma<1$;
  3. $(c)$
    $\alpha(\cdot)$ is $\mathbb{F}$-adapted and $\int^{T}_{0}\left|\alpha_{s}\right|^{2}ds<\infty$, $\mathbb{P}$-a.s.;
The set of all admissible investment strategies is denoted as $\mathcal{A}$} and is naturally defined by
[…]
We want to solve the Merton power utility optimization problem, i.e. our aim is to find the value function $\mathcal{V}(x_{0},V_{0})$ for the CRRA utility function such that
$\mathcal{V}(x_{0},V_{0})=\sup_{\alpha(\cdot)\in{\mathcal{A}}}\mathbb{E}_{x_{0},V_{0}}[\frac{1}{\gamma}(X_{T}^{\alpha})^{\gamma}];\quad 0<\gamma<1,$ (3.33)
where $\mathbb{E}_{x_{0},V_{0}}$ is the conditional expectation given $x_{0}$ and $V_{0}$. The parameter $\gamma$ represents the relative risk aversion of the investor. Smaller $\gamma$ correspond to higher risk aversion. A strategy $\alpha^{*}$ for which the supremum is attained is called an optimal strategy.
3.1.1 The degenerate correlation case
We assume that the correlation in (2.15) is of the form $(\rho,\dots,\rho)$ for $\rho\in[-1,1]$. To construct the family of processes $\{J^{\alpha}_{t}\}_{t\in[0,T]}$, $\alpha\in{\mathcal{A}}$, satisfying conditions $(1)$- $(2)$- $(3)$ in Definition 3.2, we introduce the new probability measure $\tilde{\mathbb{P}}$ defined via the Radon-Nikodym density at $\mathcal{F}_{T}$ from
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}|_{\mathcal{F}_{t}}=\mathcal{E}\Big(\frac{\gamma}{1-\gamma}\int_{0}^{t}\sum_{i=1}^{d}\theta_{i}\sqrt{V^{i}_{s}}dB^{i}_{s}\Big)=\operatorname{exp}\Big(\frac{\gamma}{1-\gamma}\int_{0}^{t}\lambda_{s}^{\top}dB_{s}-\frac{\gamma^{2}}{2(1-\gamma)^{2}}\int_{0}^{t}\left|\lambda_{s}\right|^{2}ds\Big)$
where the stochastic exponential is a true martingale by Lemma 5.1 together with the new standard brownian motion under $\tilde{\mathbb{P}}$, $\widetilde{B}_{t}=B_{t}-\frac{\gamma}{1-\gamma}\int_{0}^{t}\lambda_{s}ds$. Define the new process $\widetilde{W}$ by
$\widetilde{W}_{t}=\Sigma\widetilde{B}_{t}+\sqrt{I-\Sigma^{\top}\Sigma}B^{\perp}_{t}=W_{t}-\frac{\gamma}{1-\gamma}\int_{0}^{t}\Sigma\lambda_{s}ds,$
Notice that, by the Girsanov theorem, $\widetilde{B}$ and $\widetilde{W}$ are standard Wiener processes under the measure $\tilde{\mathbb{P}}$. As in the one dimensional case, the Ansatz ($\delta$ is called the distorsion coefficient)
$J_{t}^{\alpha}=\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Big(\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\Big(\int_{t}^{T}\frac{\gamma}{\delta}\Big(r_{s}+\frac{\left|\lambda_{s}\right|^{2}}{2(1-\gamma)}\Big)ds\Big)|\mathcal{F}_{t}\Big]\Big)^{\delta}=:\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Gamma_{t}\quad\text{where}\quad\delta:=\frac{1-\gamma}{1-\gamma+\gamma\rho^{2}}.$
is inspired by the martingale distortion transformation which was first introduced in the seminal paper Zariphopoulou (2001) and later transferred to a non-Markovian setting in Tehranchi (2004); Fouque and Hu (2019). Here we use the short notation $J_{t}^{\alpha}$ for $J_{t}(X_{t}^{\alpha},V_{t})$.
[…]
Proposition 3.4.
Assume that there exists a solution $\psi\in C([0,T],\mathbb{R}^{d})$ to the inhomogeneous Riccati-Volterra equation (3.34)-(3.35) below:
$\psi^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(\frac{\gamma\theta_{i}^{2}}{2\delta(1-\gamma)}+F_{i}(T-s,\psi(s))\big)ds,$ (3.34)
$F_{i}(s,\psi)$ $=\frac{\gamma}{1-\gamma}\rho\theta_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}+(D^{\top}\psi)_{i}+\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i})^{2},\quad i=1,\ldots,d,$ (3.35)
Then,
$\Gamma_{t}=\exp\Big(\gamma\int_{t}^{T}r(s)ds+\sum_{i=1}^{d}\int_{t}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+\delta F_{i}(s,\psi(T-s))\big)\tilde{g}^{i}_{t}(s)ds\Big).$ (3.36)
[…]
$\frac{d\Gamma_{t}}{\Gamma_{t}}$ $=\Big(-\gamma r(t)-\frac{\gamma}{2(1-\gamma)}\sum_{i=1}^{d}\Big(\theta_{i}^{2}+\delta^{2}\rho^{2}\nu_{i}^{2}(\varsigma^{i}(t)\psi^{i}(T-t))^{2}\Big)V^{i}_{t}\Big)dt+\;\delta\sum_{i=1}^{d}\psi^{i}(T-t)\nu_{i}\varsigma^{i}(t)\sqrt{V^{i}_{t}}d\widetilde{W}^{i}_{t}$
$=\big(-\gamma r(t)-\frac{\gamma}{2(1-\gamma)}\left|\lambda_{t}\right|^{2}-\frac{\gamma}{2(1-\gamma)}\delta^{2}\left|\Sigma\Lambda_{t}\right|^{2}\big)dt+\delta\Lambda_{t}^{\top}d\widetilde{W}_{t},\quad\Gamma_{T}=1.$ (3.37)
Note that Equation (5.1) is known in the litterature as a Ricatti backward stochastic differential equation (see e.g. (Abi Jaber et al., 2021, Theorem 3.1) or (Chiu and Wong, 2014, heorem 3.1), upon setting $\tilde{\Lambda}_{t}=\Gamma_{t}\Lambda_{t}$ ).
Existence and uniqueness of the solution to (3.34)-(3.35) are established in (Gnabeyeu, 2026, Theorem A.1) based on the results of Gnabeyeu et al. (b) (see also the Remark on Proposition 3.6).
By considering $\Gamma_{t}$, we overcome the non-Markovian and non-semimartingale difficulty in the variance process (2.2). Now we are ready to give the Ansatz for $J^{\alpha}_{t}$. Consider
$J^{\alpha}_{t}=\frac{X^{\gamma}_{t}}{\gamma}\Gamma_{t}.$ (3.38)
The main result we provide for this case is the following:
Theorem 3.5.
Let $\psi$ be the unique, continuous non-continuable solution of the inhomogeneous Ricatti-Volterra equation (3.34)- (3.35) on the interval $0,T_{\text{max}})$ so that Assumption 3.1 is in force. Then $J^{\alpha}_{t}=\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Gamma_{t}$ satisfies the martingale optimality principle for $t\in[0,T]$, $T<T_{\text{max}}$, and the optimal portfolio strategy $\alpha^{*}$ is given by
$\alpha^{*}_{t}$ $=\frac{1}{1-\gamma}\left(\lambda_{t}+\delta\Sigma\Lambda_{t}\right)=\frac{1}{1-\gamma}\sqrt{\operatorname{diag}(V_{t})}\left(\theta+\delta\Sigma\nu\varsigma(t)\psi(T-t)\right),\quad 0\leq t\leq T$ (3.39)
$=\;\Big(\frac{1}{1-\gamma}\big(\theta_{i}+\delta\rho\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\quad 0\leq t\leq T.$ (3.40)
Moreover, $X^{\alpha^{*}}$ satisfies (3.32) and $\alpha^{*}$ is admissible.
3.1.2 The general correlation case: A verification argument
The martingale distortion approach, used by Han and Wong (2020) in the one dimensional case can only be applied to the multivariate setting if the correlation structure is highly degenerate, i.e. $\rho_{1}=\ldots=\rho_{d}.$ (see also Fouque and Hu (2019); Aichinger and Desmettre (2021)). For the case where the correlation in (2.15) is given by an arbitrary vector $(\rho_{1},\dots,\rho_{d})\in[-1,1]^{d}$, the martingale distortion arguments from the previous section do not work anymore. However, we remark that when $\rho_{1}=\ldots=\rho_{d}=\rho$, if $\psi$ is the unique global solution of the Riccati–Volterra equation (3.34)– (3.35), then setting $\tilde{\psi}=\delta\,\psi$ and $\theta=\operatorname{diag}{(\theta_{1},\dots,\theta_{d})}$, we have using $\delta=\frac{1-\gamma}{1-\gamma+\gamma\rho^{2}}$
$\delta F(s,\psi)$ $=\frac{\gamma}{1-\gamma}\nu\varsigma(s)\Sigma\theta(\delta\psi)+D^{\top}(\delta\psi)+\frac{1}{2\delta}\Big(\nu\varsigma(s)(\delta\psi)\Big)^{2}$
[…]
where $\tilde{F}(s,\tilde{\psi}):=\frac{\gamma}{1-\gamma}\nu\varsigma(s)\Sigma\theta\tilde{\psi}+D^{\top}\tilde{\psi}+\frac{1}{2}\left[(\nu\varsigma(s)\tilde{\psi})^{2}+\frac{\gamma}{1-\gamma}(\nu\varsigma(s)\Sigma\tilde{\psi})^{2}\right]$ so that $\tilde{\psi}\in C([0,T],\mathbb{R}^{d})$ is the unique global solution of the inhomogeneous Riccati–Volterra equation
$\tilde{\psi}^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+\tilde{F}_{i}(T-s,\tilde{\psi}(s))\big)ds,$ (3.41)
[…]
Consequently, to avoid restrictions on the correlation structure linked to the martingale distortion transformation, we use a verification arguments in the spirit of Bäuerle and Li (2013) to solve the optimization problem, thus extending the results obtained in the preveous section to the more general correlation structure.
Proposition 3.6.
Assume that there exists a solution $\psi\in C([0,T],\mathbb{R}^{d})$ to the inhomogeneous Riccati-Volterra equation:
$\qquad\;\psi^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(T-s,\psi(s))\big)ds,$ (3.43)
$F_{i}(s,\psi)$ $=\frac{\gamma}{1-\gamma}\theta_{i}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}+(D^{\top}\psi)_{i}+\frac{\nu_{i}^{2}}{2}\left[(\varsigma^{i}(s)\psi^{i})^{2}+\frac{\gamma}{1-\gamma}\rho_{i}^{2}(\varsigma^{i}(s)\psi^{i})^{2}\right],\;i=1,\ldots,d.$ (3.44)
Let $\left(\Gamma,\Lambda\right)$ be defined as
$\left\{\begin{array}[]{ccl}\Gamma_{t}&=&\exp\Big(\gamma\int_{t}^{T}r(s)ds+\sum_{i=1}^{d}\int_{t}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{t}(s)ds\Big),\\ \Lambda_{t}^{i}&=&\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\sqrt{V^{i}_{t}},\quad i=1,\ldots,d,\quad 0\leq t\leq T,\end{array}\right$. (3.45)
where $g$ == $(g^{1},\ldots,g^{d})^{\top}$ is given by (2.19) i.e. the $\mathbb{R}^{d}$-valued process $(g_{t}(s))_{t\leq s}$ is defined in (2.19). Then, for some $p>1\wedge\frac{1-\gamma}{\gamma}$, $\left(\Gamma,\Lambda\right)$ is a $\mathbb{S}^{p}_{\mathbb{F}}([0,T],\mathbb{R}_{+}^{*})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$-valued solution to the Riccati BSDE (3.46)
$\left\{\begin{array}[]{ccl}\frac{d\Gamma_{t}}{\Gamma_{t}}&=&\big(-\gamma r(t)-\frac{\gamma}{2(1-\gamma)}\left|\lambda_{t}+\Sigma\Lambda_{t}\right|^{2}\big)dt+\Lambda_{t}^{\top}dW_{t},\\ \Gamma_{T}&=&1.\end{array}\right$. (3.46)
Proof: The proof that $\left(\Gamma,\Lambda\right)$ satisfy (3.46) is a straightforward adaptation of the arguments from the proof of Proposition 3.4. It remains to show that $\Gamma_{t}>0$ for all $t\in[0,T]$, $\mathbb{P}$-a.s. and $\left(\Gamma,\Lambda\right)\in\mathbb{S}^{p}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ for some sufficiently large $p>1$. An application of Itô’s formula to the process
$\bar{M}_{t}:=\Gamma_{t}\exp\Big(\int_{0}^{t}\big(\gamma r(s)+\frac{\gamma}{2(1-\gamma)}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}\big)ds\Big),\;t\leq T,$
combined with the dynamics (3.46) shows that $d\bar{M}_{t}=\bar{M}_{t}\Lambda_{t}^{\top}dW_{t}$, so that $\bar{M}$ is a local martingale of the form $\bar{M}_{t}=\;\mathcal{E}\Big(\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big)=\;\mathcal{E}\Big(\int_{0}^{t}\sum_{i=1}^{d}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V^{i}_{s}}dW^{i}_{s}\Big).$ Since $\psi$ is continuous, it is bounded; likewise, $\varsigma$ is bounded. Therefore a straightforward application of Lemma 5.1 with $g_{2}=0$ and $g_{1,i}(s)=\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\in L^{\infty}([0,T],\mathbb{R})$, recall (2.17), yields that the stochastic exponential $\bar{M}$ is a true $\mathbb{P}$- martingale. Now, as $\Gamma_{T}=1$, writing $\mathbb{E}[\bar{M}_{T}|\mathcal{F}_{t}]=\bar{M}_{t}$, we obtain
$\Gamma_{t}$ $=\mathbb{E}\Big[\exp\Big(\int_{t}^{T}\big(\gamma r(s)+\frac{\gamma}{2(1-\gamma)}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}\big)ds\Big)\mid\mathcal{F}_{t}\Big],\quad t\leq T$ (3.47)
$\leq\exp\Big(\int_{0}^{T}\gamma r(s)ds\Big)M_{t},\;\text{with}\;M_{t}:=\mathbb{E}\Big[\exp\Big(\frac{\gamma}{2(1-\gamma)}\int_{0}^{T}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}ds\Big)\mid\mathcal{F}_{t}\Big],\quad t\leq T$
We then have in view of (3.47) that there exists some positive constant $m>0$ such that $\Gamma_{t}\geq m>0$ for every $t\in[0,T]$. Using the elementary inequality $(a+b)^{p}\leq 2^{(p-1)^{+}}\big(a^{p}+b^{p}\big)$ for $a,b>0$, one gets
$|\lambda_{s}+\Sigma\Lambda_{s}|^{2}\leq 2(|\lambda_{s}|^{2}+|\Sigma\Lambda_{s}|^{2})\leq 2(1+|\Sigma|^{2})(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2})\leq 2\beta(1+|\Sigma|^{2})(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}).$
where $\beta>1$ is choosen such that $\frac{\beta\gamma}{(1-\gamma)}>1$. Consequently, by simple calculation:
$\mathbb{E}\Big[\exp\Big(\frac{\gamma}{2(1-\gamma)}\int_{0}^{T}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}ds\Big)\Big]$ $\leq\mathbb{E}\Big[\exp\Big(\frac{\beta\gamma}{(1-\gamma)}(1+|\Sigma|^{2})\int_{0}^{T}\big(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\big)ds\Big)\Big]<\infty$
thanks to the Novikov-type condition (3.27) with constant $a(\frac{\beta\gamma}{(1-\gamma)})=\frac{\beta\gamma}{(1-\gamma)}(1+|\Sigma|^{2})$. Therefore, $M_{t}$ is a martingale under $\mathbb{P}$. Let $p>1\wedge\frac{1-\gamma}{\gamma}$. By virtue of Doob’s maximal inequality, in the second line,
$\,\mathbb{E}\Big[\sup_{t\in[0,T]}\big|\Gamma_{t}\big|^{p}\Big]\leq e^{p\gamma\int_{0}^{T}r(s)ds}\mathbb{E}\Big[\sup_{t\in[0,T]}\big|M_{t}\big|^{p}\Big]$
$\hskip 71.13188pt\leq e^{p\gamma\int_{0}^{T}r(s)ds}\Big(\frac{p}{p-1}\Big)^{p}\mathbb{E}\Big[\exp\Big(\frac{p\gamma}{2(1-\gamma)}\int_{0}^{T}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}ds\Big)\Big]$
$\hskip 71.13188pt\leq\Big(\frac{p}{p-1}\Big)^{p}e^{p\gamma\int_{0}^{T}r(s)ds}\mathbb{E}\left[\exp\left({a(\frac{p\gamma}{(1-\gamma)})\int_{0}^{T}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds}\right)\right]<\infty$
where we used condition (3.27) with $a(\frac{p\gamma}{(1-\gamma)})=\frac{p\gamma}{(1-\gamma)}(1+|\Sigma|^{2})$. Therefore, $\mathbb{E}\big[\sup_{t\in[0,T]}|\Gamma_{t}|^{p}\big]<\infty$ holds. As for $\Lambda$, it is clear that it belongs to $L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ since $\varsigma$ and $\psi$ are bounded, $\psi$ is continuous thus bounded and $\mathbb{E}\Big[\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\Big]<\infty$ by (2.17). This complete the Proof $\Box$
First of all, we point out that for each of the inhomogeneous Riccati-Volterra equations (3.43)- (3.44) and (3.34)- (3.35), there exists a unique non-continuable continuous solution $(\psi,T_{\operatorname{max}})$ under Assumption 2.1 by (Gnabeyeu, 2026, Theorem A.1) as stated in the following Remark.
Remark: Fix $T>0$ and assume that $K$ satisfies Assumption 2.1. As the matrix $D$ in the drift of the volatility process is a diagonal matrix, i.e. $D=-\operatorname{diag}{(\lambda_{1},\dots,\lambda_{d})}$, for $i=1,\ldots,d,$ $\psi^{i}$ satisfies the Volterra equation
$\chi(t)=\int_{0}^{t}K_{i}(t-s)\Big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}-\big(\lambda_{i}-\frac{\gamma}{1-\gamma}\theta_{i}\rho_{i}\nu_{i}\varsigma^{i}(T-s)\big)\chi(s)+\frac{\nu_{i}^{2}}{2}\frac{1-\gamma+\gamma\rho_{i}^{2}}{1-\gamma}\varsigma^{i}(T-s)^{2}\chi(s)^{2}\Big)ds,\;t\leq T.$ (3.48)
(Gnabeyeu, 2026, Theorem A.1) guarantees that there exists a unique non-continuable continuous solution $(\psi,T_{\max})$ to Equation (3.74) with $\psi^{i}\in C([0,T_{max}),\mathbb{R}^{d})$ in the sense that $\psi^{i}$ satisfies (3.74) on $0,T_{max})$ with $T_{max}\in(0,T]$ and $\sup_{t<T_{max}}|\psi^{i}(t)|=+\infty$, if $T_{max}<T$. Combining the component-wise solutions, we finally obtain the existence of a unique non-continuable solution $\psi\in C([0,T_{max}),\mathbb{R}^{d})$ of the inhomogeneous Ricatti–Volterra Equation (3.43)- (3.44) (and in particular of (3.34)– (3.35) ).
Note that our candidate for the optimal portfolio strategy $\alpha^{*}$ is given, using (3.39)– (3.40), by
$\alpha_{t}^{*}=\frac{1}{1-\gamma}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)=\;\Big(\frac{1}{1-\gamma}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\;0\leq t\leq T.$ (3.49)
We will show directly that the corresponding value attained by this portfolio strategy is $\frac{x_{0}^{\gamma}}{\gamma}\Gamma_{0}$ and every other admissible portfolio strategy will not yield a larger value. Then we have the following verification result:
Theorem 3.7.
Let $\psi$ be the unique, continuous non-continuable solution of the inhomogeneous Ricatti-Volterra equation (3.43)- (3.44) on the interval $0,T_{\text{max}})$ so that Assumption 3.1 is in force. Then for $t\in[0,T]$, $T<T_{\operatorname{max}}$, an optimal investment strategy $(\alpha_{t}^{*})_{t\in[0,T]}$ for the Merton portfolio problem (3.33) is given by
$\alpha^{*}_{t}$ $=\frac{1}{1-\gamma}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)=\frac{1}{1-\gamma}\sqrt{\operatorname{diag}(V_{t})}\left(\theta+\Sigma\nu\varsigma(t)\psi(T-t)\right),\quad 0\leq t\leq T$ (3.50)
$=\;\Big(\frac{1}{1-\gamma}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\quad 0\leq t\leq T.$ (3.51)
Moreover,
$\mathbb{E}\Big[\sup_{t\in[0,T]}|X^{\alpha^{*}}_{t}|^{p}\Big]<\infty,\quad\text{for some sufficiently large}\quad p>1.$ (3.52)
and $\alpha^{*}$ is admissible. The value function defined in (3.33) can be written as
$\sup_{\alpha\in\mathcal{A}}\mathbb{E}_{x_{0},V_{0}}\left[\frac{\big(X_{T}^{\alpha^{*}}\big)^{\gamma}}{\gamma}\right]=\mathcal{V}(x_{0},V_{0})=\frac{x_{0}^{\gamma}}{\gamma}\exp\Big(\gamma\int_{0}^{T}r(s)ds+\sum_{i=1}^{d}\int_{0}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$
For the sake of space limitation, but without loss of self-containedness, we present a sketch of the proof here.
Sketch of Proof: In order to prove that $\alpha^{*}$ is indeed the optimal portfolio strategy, we show that for
$G(x_{0},V_{0}):=\frac{x_{0}^{\gamma}}{\gamma}\Gamma_{0}=\frac{x_{0}^{\gamma}}{\gamma}\exp\Big(\gamma\int_{0}^{T}r(s)ds+\sum_{i=1}^{d}\int_{0}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big),$
we have
  1. 1.
    $\mathbb{E}_{x_{0},V_{0}}{\big[\frac{(X_{T}^{\alpha^{*}})^{\gamma}}{\gamma}\big]}=G(x_{0},V_{0})$ for $\alpha_{t}^{*}=\frac{\lambda_{t}+\Sigma\Lambda_{t}}{1-\gamma}=\Big(\frac{1}{1-\gamma}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\;t\in[0,T]$,
  2. 2.
    $\mathbb{E}_{x_{0},V_{0}}{\big[\frac{(X_{T}^{\alpha})^{\gamma}}{\gamma}\big]}\leq G(x_{0},V_{0})$ for every other admissible strategy $\alpha\in\mathcal{A}$.
In fact, inserting the candidate $\alpha^{*}$ for the optimal portfolio strategy in the SDE of the wealth equation show by $(1)$ that the upper bound $G$ can be obtained. Consequently conditions $(1)$ and $(2)$ ensure that $G$ is the value function of the problem (3.33) and $\alpha^{*}$ is the optimal portfolio strategy.
3.2 Optimal strategy for the exponential utility maximization problem
In this subsection, we consider the exponential utility case. With a slightly different formulation, let $\pi_{t}$ denote the vector of the amounts invested in the risky assets $S$ at time $t$ in a self–financing strategy. We assume that the the process $(\pi_{t})_{t\geq 0}$ are progressively measurable. Then, the dynamics of the wealth $X^{\pi}$ of the portfolio we seek to optimize is given by
$\mathrm{d}X^{\pi}_{t}$ $=\bigl(\pi_{t}^{\top}\big(r(t){\mathbb{1}_{d}}+\sigma(V_{t})\lambda_{t}\big)+(X^{\pi}_{t}-\pi_{t}^{\top}{\mathbb{1}_{d}})r(t)\bigr)\,\mathrm{d}t+\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}$
$=X^{\pi}_{t}\bigl(r(t)+\pi_{t}^{\top}\sigma(V_{t})\lambda_{t}\bigr)\,\mathrm{d}t+\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}=X_{t}^{\pi}(r(t)+\pi_{t}^{\top}{\operatorname{diag}(V_{t})}{\theta})dt+\pi_{t}^{\top}\sqrt{\operatorname{diag}(V_{t})}dB_{t},$
Letting $\alpha_{t}:=\sigma^{\top}(V_{t})\pi_{t}$ be the investment strategy, the wealth $X^{\alpha}$ reads:
$dX^{\alpha}_{t}$ $=\big(r(t)X^{\alpha}_{t}+\alpha_{t}^{\top}\lambda_{t}\big)dt+\alpha_{t}^{\top}dB_{t},\quad t\geq 0,\quad X_{0}^{\alpha}=x_{0}\in\mathbb{R}.$ (3.53)
By a standard calculation, the wealth process is then given by
$X_{t}=e^{\int^{t}_{0}r(s)ds}\left(x_{0}+\int_{0}^{t}e^{-\int^{s}_{0}r(u)du}\left(\alpha_{s}^{\top}dB_{s}+\alpha_{s}^{\top}\lambda_{s}\mathrm{d}s\right)\right),$ (3.54)
Note that it is sufficient to assume that $\int_{0}^{t}(\left|\lambda_{s}\right|^{2}+\left|\alpha_{s}\right|^{2})\,\mathrm{d}s<+\infty$ almost surely for all $t\geq 0$ in order to construct the stochastic integrals in Equation (3.54). This boundedness condition holds owing to the inequality $|z|^{2}\leq 2e^{|z|},\,\forall z\in\mathbb{R}$, together with the condition (3.27), and the following admissibility assumption, which is consistent with Hu et al. (2005).
Definition 3.8.
Let $\gamma>0$, an investment strategy $\alpha(\cdot)$ is said to be admissible if
  1. $(a)$
    $\alpha(\cdot)$ is $\mathbb{F}$-adapted and $\mathbb{E}[\int^{T}_{0}\left|\alpha_{s}\right|^{2}ds]<\infty$;
  2. $(b)$
    The wealth process (3.53) has a unique solution in terms of $(S,V,B)$, with $\mathbb{P}$-a.s. continuous paths;
  3. $(c)$
    $\left\{\exp\big[-\gamma e^{\int^{T}_{\tau}r(u)du}X_{\tau}\big]:\tau\text{ stopping time with values in }[0,T]\right\}$ is a uniformly integrable family.
In particular,
$\sup_{\tau\in[0,T]}\mathbb{E}\Big[\exp\left(-p\gamma e^{\int^{T}_{\tau}r(u)du}X^{\alpha}_{\tau}\right)\Big]<\infty,\text{ for some }p>1.$ (3.55)
The set of all admissible investment strategies is denoted as $\mathcal{A}$} and is naturally defined by
[…]
The investor now considers the Merton exponential utility optimization problem, i.e. its aim is to find the value function $\mathcal{V}(x_{0},V_{0})$ for the CARA utility function such that
$\mathcal{V}(x_{0},V_{0})=\sup_{\alpha(\cdot)\in{\mathcal{A}}}\mathbb{E}_{x_{0},V_{0}}\left[-\frac{1}{\gamma}\exp\Big(-\gamma X_{T}\Big)\right],\quad\gamma>0.$ (3.56)
The solution method is similar to the power utility case. With slightly abuse of notations, we still use $\psi(\cdot)$, $\Gamma$, etc. However, they are redefined and not mixed with counterparts in power utility case.
3.2.1 The degenerate correlation case
We assume that the correlation in (2.15) is of the form $(\rho,\dots,\rho)$ for $\rho\in[-1,1]$. To construct the family of processes $\{J^{\alpha}_{t}\}_{t\in[0,T]}$, $\alpha\in{\mathcal{A}}$, satisfying conditions $(1)$- $(2)$- $(3)$ in Definition 3.2, we introduce the new probability measure $\tilde{\mathbb{P}}$ defined via the Radon-Nikodym density or derivative at $\mathcal{F}_{T}$ from
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}|_{\mathcal{F}_{t}}=\mathcal{E}\Big(-\int_{0}^{t}\sum_{i=1}^{d}\theta_{i}\sqrt{V^{i}_{s}}dB^{i}_{s}\Big)=\operatorname{exp}\Big(-\int_{0}^{t}\lambda_{s}^{\top}dB_{s}-\frac{1}{2}\int_{0}^{t}\left|\lambda_{s}\right|^{2}ds\Big)$
where the stochastic exponential is a true martingale by Lemma 5.1 together with the new standard brownian motion under $\tilde{\mathbb{P}}$ , $\widetilde{B}_{t}=B_{t}+\int_{0}^{t}\lambda_{s}ds.$ Define the new process $\widetilde{W}$ by
$\widetilde{W}_{t}=\Sigma\widetilde{B}_{t}+\sqrt{I-\Sigma^{\top}\Sigma}B^{\perp}_{t}=W_{t}+\int_{0}^{t}\Sigma\lambda_{s}ds,$
Notice that, by the Girsanov theorem, $\widetilde{B}$ and $\widetilde{W}$ are standard Wiener processes under the measure $\tilde{\mathbb{P}}$. As in the one dimensional case (see e.g. Han and Wong (2020)), the Ansatz
$J_{t}^{\alpha}=-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\big)\Big(\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\left|\lambda_{s}\right|^{2}ds\big)|\mathcal{F}_{t}\Big]\Big)^{\frac{1}{1-\rho^{2}}}=:-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\big)\Gamma_{t}.$
is inspired by the martingale distortion transformation in a non-Markovian setting in Tehranchi (2004) for the case of the exponential utility function, where the distortion power arises from simple Hölder-type inequalities. Here we still use the short notation $J_{t}^{\alpha}$ for $J_{t}(X_{t}^{\alpha},V_{t})$.
Proposition 3.9.
Assume that there exists a solution $\psi\in C([0,T],\mathbb{R}^{d})$ to the inhomogeneous Riccati-Volterra equation (3.57)-(3.58) below:
$\psi^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(T-s,\psi(s))\big)ds,$ (3.57)
$F_{i}(s,\psi)$ $=-\rho\theta_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}+(D^{\top}\psi)_{i}+\frac{\nu_{i}^{2}}{2}(1-\rho^{2})(\varsigma^{i}(s)\psi^{i})^{2},\quad i=1,\ldots,d,$ (3.58)
Then,
$\Gamma_{t}=\exp\Big(\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)\tilde{g}^{i}_{t}(s)ds\Big).$ (3.59)
[…]
$\Lambda_{t}^{i}=\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\sqrt{V^{i}_{t}},\quad i=1,\ldots,d,\quad 0\leq t\leq T,$ (3.60)
Moreover, $\left(\Gamma,\Lambda\right)\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ and
$d\Gamma_{t}$ $=\;\Gamma_{t}\Big(\sum_{i=1}^{d}\Big(\frac{\theta_{i}^{2}}{2}+\frac{\nu_{i}^{2}}{2}\rho^{2}(\varsigma^{i}(t)\psi^{i}(T-t))^{2}\Big)V^{i}_{t}+\sum_{i=1}^{d}\psi^{i}(T-t)\nu_{i}\varsigma^{i}(t)\sqrt{V^{i}_{t}}d\widetilde{W}^{i}_{t}\Big)$ (3.61)
$=\;\Gamma_{t}\Big[\frac{1}{2}\big(\left|\lambda_{t}\right|^{2}+\left|\Sigma\Lambda_{t}\right|^{2}\big)dt+\Lambda_{t}^{\top}d\widetilde{W}_{t}\Big],$ (3.62)
Remark Assume that $K$ satisfies the Assumption 2.1. As $1-\rho^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.1) provides the existence of a unique global continuous solution on $[0,T]$ to (3.57)– (3.58).
Consider the below Ansatz for $J^{\alpha}_{t}$
$J^{\alpha}_{t}=-\frac{1}{\gamma}\exp\big[-\gamma e^{\int^{T}_{t}r(s)ds}X_{t}^{\alpha}\big]\Gamma_{t}.$ (3.63)
The main result we provide for this case is the following:
Theorem 3.10.
Let $\psi$ be the unique, continuous solution of the inhomogeneous Ricatti-Volterra equation (3.57)- (3.58) on the interval $[0,T]$, such that Assumption 3.1 is in force. Then $J^{\alpha}_{t}$ in (3.63) satisfies the martingale optimality principle for $t\in[0,T]$, and the optimal portfolio strategy $\alpha^{*}$ is given by
$\alpha^{*}_{t}$ $=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\sqrt{\operatorname{diag}(V_{t})}\left(\theta+\Sigma\nu\varsigma(t)\psi(T-t)\right),\quad 0\leq t\leq T$ (3.64)
$=\;\Big(\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\quad 0\leq t\leq T.$ (3.65)
Moreover, $X^{\alpha^{*}}$ satisfies Equation (3.55) and $\alpha^{*}$ is admissible.
3.2.2 The general correlation case: A verification argument
As developed in the previous section, when the correlation structure is highly degenerate, that is, when $\rho_{1}=\ldots=\rho_{d}$, the martingale distortion approach used by Han and Wong (2020) in the one-dimensional case extends naturally to the multivariate setting. When the correlation structure in (2.15) is given by an arbitrary vector $(\rho_{1},\dots,\rho_{d})\in[-1,1]^{d}$, we rewrite the Riccati–Volterra equations (3.57)–(3.58) as
$\psi^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\Big(-\frac{\theta_{i}^{2}}{2}+F_{i}(T-s,\psi(s))\Big)\,ds,$ (3.66)
$F_{i}(s,\psi)$ $=-\theta_{i}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}+(D^{\top}\psi)_{i}+\frac{\nu_{i}^{2}}{2}(1-\rho_{i}^{2})\big(\varsigma^{i}(s)\psi^{i}\big)^{2},\quad i=1,\ldots,d.$ (3.67)
First, note that if $K_{i}$ satisfies the Assumption 2.1 for $i=1,\ldots,d$. As $1-\rho_{i}^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.1) provides the existence of a unique global continuous solution on $[0,T]$ to (3.66)– (3.67). More details are given in the Remark on Proposition 3.11. In particular, when $\rho_{1}=\ldots=\rho_{d}=\rho$, the function $\psi$ coincides with the unique global solution to the Riccati–Volterra equations (3.57)–(3.58). Consequently, to avoid restrictions on the correlation structure, we still use a verification arguments in the spirit of Bäuerle and Li (2013) to solve the optimization problem, thus extending the results obtained in the preveous section to the more general correlation structure.
Proposition 3.11.
Assume that there exists a solution $\psi\in C([0,T],\mathbb{R}^{d})$ to the inhomogeneous Riccati-Volterra equation:
$\psi^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(T-s,\psi(s))\big)ds,$ (3.68)
$F_{i}(s,\psi)$ $=-\theta_{i}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}+(D^{\top}\psi)_{i}+\frac{\nu_{i}^{2}}{2}(1-\rho^{2}_{i})(\varsigma^{i}(s)\psi^{i})^{2},\quad i=1,\ldots,d,$ (3.69)
Let $\left(\Gamma,\Lambda\right)$ be defined as
$\left\{\begin{array}[]{ccl}\Gamma_{t}&=&\exp\Big(\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{t}(s)ds\Big),\\ \Lambda_{t}^{i}&=&\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\sqrt{V^{i}_{t}},\quad i=1,\ldots,d,\quad 0\leq t\leq T,\end{array}\right$. (3.70)
where $g$ == $(g^{1},\ldots,g^{d})^{\top}$ is given by (2.19) i.e. the $\mathbb{R}^{d}$-valued process $(g_{t}(s))_{t\leq s}$ is defined in (2.19). Then, $\left(\Gamma,\Lambda\right)$ is a $\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$-valued solution to the Riccati BSDE (3.71) below.
$\left\{\begin{array}[]{ccl}\frac{d\Gamma_{t}}{\Gamma_{t}}&=&\frac{1}{2}\left|\lambda_{t}+\Sigma\Lambda_{t}\right|^{2}\,dt+\Lambda_{t}^{\top}dW_{t},\\ \Gamma_{T}&=&1.\end{array}\right$. (3.71)
Moreover, $\Gamma_{t}$ is essentially bounded. Specifically, $0<\Gamma_{t}\leq 1$ for all $t\in[0,T]$, $\mathbb{P}$-a.s..
Proof: The proof that $\left(\Gamma,\Lambda\right)$ satisfy (3.71) is a straightforward adaptation of the arguments from the proof of Proposition 3.9. It remains to show that $0<\Gamma_{t}\leq 1$ for all $t\in[0,T]$, $\mathbb{P}$-a.s. and $\left(\Gamma,\Lambda\right)\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. For this, define the process
$M_{t}$ $=\Gamma_{t}\exp\Big(-\frac{1}{2}\int_{0}^{t}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}ds\Big)=\Gamma_{t}\exp\Big(-\frac{1}{2}\int_{0}^{t}\sum_{i=1}^{d}V^{i}_{s}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\big)^{2}ds\Big),\quad t\leq T.$
An application of Itô’s formula combined with the dynamics (3.71) shows that $dM_{t}=M_{t}\Lambda_{t}^{\top}dW_{t}$, so that $M$ is a local martingale of the form
$M_{t}$ $=\;\mathcal{E}\Big(\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big)=\;\mathcal{E}\Big(\int_{0}^{t}\sum_{i=1}^{d}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V^{i}_{s}}dW^{i}_{s}\Big).$ (3.72)
Since $\psi$ is continuous, it is bounded; likewise, $\varsigma$ is bounded. Therefore a straightforward application of Lemma 5.1 with $g_{2}=0$ and $g_{1,i}(s)=\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\in L^{\infty}(\mathbb{R}^{+},\mathbb{R})$, recall (2.17), yields that the stochastic exponential $M$ is a true $\mathbb{P}$- martingale. Now, as $\Gamma_{T}=1$, writing $\mathbb{E}[M_{T}|\mathcal{F}_{t}]=M_{t}$, we obtain
$\Gamma_{t}=\mathbb{E}\Big[\exp\Big(-\frac{1}{2}\int_{t}^{T}\sum_{i=1}^{d}V^{i}_{s}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\big)^{2}ds\Big)\mid\mathcal{F}_{t}\Big],\quad t\leq T,$ (3.73)
which ensures that $0<\Gamma_{t}\leq 1$, $\mathbb{P}-a.s.$, since $V_{t}$ is non-negative ($V\in\mathbb{R}^{d}_{+}$). As for $\Lambda$, it is clear that it belongs to $L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ since $\varsigma$ and $\psi$ are bounded, $\psi$ is continuous thus bounded and $\mathbb{E}\Big[\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\Big]<\infty$ by (2.17). This complete the Proof $\Box$
Remark: Assume that $K$ satisfies the Assumption 2.1. Then (Gnabeyeu, 2026, Theorem A.1) provides the existence of a unique global continuous solution $\psi\in C([0,T],\mathbb{R}^{d})$ to (3.68)– (3.69) (and in particular to (3.57)– (3.58) ) and $\psi<0$ for $t>0$. More precisely, as the matrix $D$ in the drift of the volatility process is a diagonal matrix, i.e. $D=-\operatorname{diag}{(\lambda_{1},\dots,\lambda_{d})}$, for $i=1,\ldots,d,$ by (Gnabeyeu, 2026, Corollary A.2), since $\frac{\theta_{i}^{2}}{2}<0$, $\psi^{i}\in C([0,T],\mathbb{R}_{-})$ is unique global solution to the following Volterra equation
$\chi(t)=\int_{0}^{t}K_{i}(t-s)\Big(-\frac{\theta_{i}^{2}}{2}-\big(\lambda_{i}+\theta_{i}\rho_{i}\nu_{i}\varsigma^{i}(T-s)\big)\chi(s)+\frac{\nu_{i}^{2}}{2}(1-\rho_{i}^{2})\varsigma^{i}(T-s)^{2}\chi(s)^{2}\Big)ds,\;t\leq T.$ (3.74)
Combining the component-wise solutions, we finally obtain the unique global solution $\psi$ of the inhomogeneous Ricatti–Volterra Equation (3.68)– (3.68) (and in particular to (3.57)– (3.58) ).
Moreover, it follows in this case that the condition (3.25) can be made more explicit by bounding $\psi$ with respect to the vector $\theta$. Indeed setting for $i=1,\ldots,d,$ $\bar{\lambda}_{i}:=\inf_{t\in[0,T]}\big(\lambda_{i}+\nu_{i}\rho_{i}\theta_{i}\varsigma^{i}(t)\big)=\lambda_{i}+\nu_{i}\rho_{i}\theta_{i}\|\varsigma^{i}\|_{\infty}{\mathbb{1}_{\rho_{i}\leq 0}},$ which is strictly positive, still by (Gnabeyeu, 2026, Corollary A.2), we have:
$\sup_{t\in[0,T]}|\psi^{i}(t)|\leq\frac{|\theta_{i}|^{2}}{2\bar{\lambda}_{i}}\int_{0}^{T}f_{\bar{\lambda}_{i}}(s)ds=\frac{|\theta_{i}|^{2}}{2\bar{\lambda}_{i}}(1-R_{\bar{\lambda}_{i}}(T)),\;i=1,\ldots,d.$ (3.75)
where $R_{\bar{\lambda}_{i}}$ is the $\bar{\lambda}_{i}$-resolvent associated to the real-valued kernel $K_{i}$ and $f_{\bar{\lambda}_{i}}$ its antiderivative. Consequently, combining those component-wise estimates, we finally obtain that, a sufficient condition on $\theta$ to ensure (3.25) would be
$\theta^{2}_{i}\left(1+(\nu_{i}\|\varsigma^{i}\|_{\infty}\frac{\theta_{i}}{2\bar{\lambda}_{i}})^{2}\big(1-R_{\bar{\lambda}_{i}}(T)\big)^{2}\right)\leq\frac{a}{a(p)}\quad\text{for all}\quad i=1,\ldots,d.$ (3.76)
Now, observe that the proposed candidate for the optimal portfolio strategy $\alpha^{*}$ follows from (3.64)–(3.65) and is given by
$\alpha_{t}^{*}=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)=\Bigg(\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Bigg)_{1\leq i\leq d},\quad 0\leq t\leq T.$ (3.77)
We will show directly that this strategy attains the value $\frac{1}{\gamma}\exp\left(-\gamma e^{\int_{t}^{T}r(s)ds}x_{0}\right)\Gamma_{0}$, and that no other admissible portfolio strategy can achieve a higher value. The main verification result for this setting is stated as follows:
Theorem 3.12.
Let $\psi$ be the unique, continuous solution of the inhomogeneous Ricatti-Volterra equation (3.68)- (3.69) on the interval $[0,T]$, such that Assumption 3.1 is in force. Then for $t\in[0,T]$, an optimal investment strategy $(\alpha_{t}^{*})_{t\in[0,T]}$ for the Merton portfolio problem (3.56) is given by
$\alpha^{*}_{t}$ $=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\sqrt{\operatorname{diag}(V_{t})}\left(\theta+\Sigma\nu\varsigma(t)\psi(T-t)\right),\quad 0\leq t\leq T$ (3.78)
$=\;\Big(\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\quad 0\leq t\leq T.$ (3.79)
Moreover,
$\sup_{\tau\in[0,T]}\mathbb{E}\Big[\exp\left(-p\gamma e^{\int^{T}_{\tau}r(u)du}X^{\alpha^{*}}_{\tau}\right)\Big]<\infty,\text{ for some }p>1.$ (3.80)
and $\alpha^{*}$ is admissible. The value function defined in (3.56) can be written as
$\mathcal{V}(x_{0},V_{0})=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$ (3.81)
For the sake of brevity, yet without compromising self-containment, we present a sketch of the proof here.
Sketch of Proof: In order to prove that $\alpha^{*}$ is indeed the optimal portfolio strategy, we show that for
$G(x_{0},V_{0}):=-\frac{\Gamma_{0}}{\gamma}\exp\big(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\big)=-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}+\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\big)$
we have
  1. 1.
    $\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\big(-\gamma X_{T}^{\alpha^{*}}\big)\big]}=G(x_{0},V_{0})$ for $\alpha_{t}^{*}=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right),\;0\leq t\leq T$,
  2. 2.
    $\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\big(-\gamma X_{T}^{\alpha}\big)\big]}\leq G(x_{0},V_{0})$ for every other admissible strategy $\alpha\in\mathcal{A}$.
In fact, conditions $(1)$ and $(2)$ ensure that G is the value function of the problem (3.56) and $\alpha^{*}$ is the optimal portfolio strategy.
4 Numerical experiments: The particular case of the fake stationary rough Heston volatility
In this section, we illustrate the results of Section 2 by numerically computing the optimal portfolio strategy for a special case of two-dimensional fake stationary rough heston model model as described in section 3. We consider a financial market consisting of one risk-free asset and $d=2$ risky assets, with an investment horizon of $T=1$ year. To model the roughness of the asset price dynamics, we employ an appropriate integration kernel. We choose a fractional kernel of Remark 2.1 and Example 2.6 of the form:
$K(t)=\begin{pmatrix}\frac{t^{\alpha_{1}-1}}{\Gamma(\alpha_{1})}&0\\ 0&\frac{t^{\alpha_{2}-1}}{\Gamma(\alpha_{2})}\end{pmatrix},\quad 0.4+\frac{1}{2}=\alpha_{1},\,0.1+\frac{1}{2}=\alpha_{2}\in\big(\frac{1}{2},1\big).$
Here, $\Gamma(\alpha)$ is the Gamma function, and the parameter $\alpha$ controls the degree of roughness in the model.
Note that, the model is sufficiently rich to capture several well-known stylized facts of financial markets:
  • Each asset $S^{i}$, $i=1,2$ exhibits stochastic rough volatility driven by the process $V^{i}$, with different Hurst indices $\alpha_{i}$.
  • Each stock $S^{i}$ is correlated with its own volatility process through the parameter $\rho_{i}$ to take into account the leverage effect.
We consider the setting where in Equation (2.2), the simplified specification $\varphi(t)=I_{2\times 2},\;t\geq 0,$ holds almost surely, in which case the $\mathbb{R}^{d}$- valued mean-reverting function $\mu$ is constant in time, that is, $\mu(t)=\mu_{0}\in\mathbb{R}^{2},\;\forall\,t\geq 0,$ (see, e.g., Gnabeyeu and Pagès (2025)).
We consider the following estimates for the model parameters $V_{0}^{i}\sim\mathcal{N}(\frac{\mu_{0}^{i}}{\lambda_{i}},v_{0}^{i})$ defined in (2.12) $($$E_{\lambda_{i},c_{i}}$)):
$c=\begin{pmatrix}0.01\\ 0.03\end{pmatrix},\quad\mu_{0}=\begin{pmatrix}0.2\\ 0.25\end{pmatrix},\;D=\begin{pmatrix}-0.2&0\\ 0&-0.6\end{pmatrix},\quad\Sigma=\begin{pmatrix}-0.7&0\\ 0&-0.55\end{pmatrix},\;\theta=\begin{pmatrix}0.1\\ 0.1\end{pmatrix},\;\nu=\begin{pmatrix}0.4\\ 0.2\end{pmatrix}$.
Remark: In order to numerically implement the optimal strategy (3.51)–(3.79), one needs to simulate the non-Markovian process $V$ in Equation (2.2)– (2.18) and to discretize the Riccati-Volterra equation for $\psi$ in (3.43)– (3.44) and (3.68)– (3.69) respectively.
4.1 About the numerical scheme for the Volterra and fractional Riccati equations
To simulate the Volterra process (2.2)–(2.18), we first rewrite it using the equivalent Wiener–Hopf transform (see, e.g., (Gnabeyeu and Pagès, 2025; Gnabeyeu et al., a, Propositions 2.8 and 2.2)). We then introduce the $f_{\lambda}$-integrated discrete time Euler-Maruyama scheme defined by the below equation on the time grid $t_{k}=t^{n}_{k}=\frac{kT}{n},k=0,\dots,n$ and in the fractional kernel case, namely recursively for $i=1,2$, $\overline{V}^{i,n}_{0}=V^{i,0}$ and for every $k=1,\ldots,n$,
$\overline{V}^{i,n}_{t_{k}}=\frac{\mu_{0}^{i}}{\lambda_{i}}+\big(V^{i,0}-\frac{\mu_{0}^{i}}{\lambda_{i}}\big)R_{\lambda_{i}}(t_{k})+\frac{\nu_{i}}{\lambda_{i}}\sum_{\ell=1}^{k}\varsigma^{i}(t_{\ell})\sqrt{\overline{V}_{t_{\ell-1}}^{i,n}}\int_{t_{\ell-1}}^{t_{\ell}}f_{\alpha_{i},\lambda_{i}}(t_{k}-s)\,dW_{s}=h^{i}(t_{k})+\frac{\nu_{i}}{\lambda_{i}}\sum_{\ell=1}^{k}\varsigma^{i}(t_{\ell})\sqrt{\overline{V}_{t_{\ell-1}}^{i,n}}I^{i,\ell}_{k}$.
where the stochastic integrals $\left(I^{i,\ell}_{k}=\int_{t_{\ell-1}}^{t_{\ell}}f_{\alpha_{i},\lambda_{i}}(t_{k}-s)\,dW_{s}\right)$ can be simulated on the discrete grid $(t^{n}_{k})_{0\leq k\leq n}$ by generating an independent sequence of gaussian vectors $G^{n,\ell},\ell=1\cdots n$ using an extended and stable version of Cholesky decomposition of a well-defined covariance matrix $C$. The reader is referred to (Gnabeyeu and Pagès, 2025, Appendix A) for further details about the simulation of the Gaussian stochastic integrals terms in the semi-integrated Euler scheme introduced in this context for Equation (2.2)– (2.18).
To design an approximation scheme for solving the two-dimensional Riccati–Volterra system of equations (3.43)–(3.44) and (3.68)–(3.69) numerically, we employ as in (El Euch and Rosenbaum, 2019, section 5.1) the generalized Adams–Bashforth–Moulton method, often referred to as the fractional Adams method investigated in Diethelm et al. (2002, 2004) as a useful numerical algorithm for solving a fractional ordinary differential equation (ODE) based on a predictor-corrector approach.
Over a regular or uniform discrete time-grid $(t_{k})_{k=1,\ldots,n}$ with mesh or step length $\Delta=\frac{T}{n}$ $(t_{k}=k\Delta)$ for some integer $n$, let $y_{j}^{i}:=\psi^{i}(t_{j})$, the explicit numerical scheme to estimate $\psi$ is given by
$\begin{cases}y_{k+1}^{i,P}=\displaystyle\sum_{j=0}^{k}b^{i}_{j,k+1}f_{i}(t_{j},y_{j}^{i}),\;\text{with}\;f_{i}(t_{j},x):=a_{i}+F_{i}(T-t_{j},x)\\ y_{k+1}^{i}=\displaystyle\sum_{j=0}^{k}a^{i}_{j,k+1}f_{i}(t_{j},y_{j}^{i})+a^{i}_{k+1,k+1}f_{i}(t_{k+1},y_{k+1}^{i,P}),\quad y_{0}^{i}=0\end{cases},\;i\in\{1,2\}.$
where $a_{i}=\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}$ in section 3.1 (resp. $a_{i}=-\frac{\theta_{i}^{2}}{2}$ in section 3.2), $F_{i}$ in (3.44) (resp. in (3.69)) and the weights $a^{i}_{j,k+1}$, $b^{i}_{j,k+1}$ are defined as
$a^{i}_{j,k+1}=\frac{\Delta^{\alpha_{i}}}{\Gamma(\alpha_{i}+2)}\begin{cases}k^{\alpha_{i}+1}-(k-\alpha_{i})(k+1)^{\alpha_{i}},&\text{if }j=0,\\ (k-j+2)^{\alpha_{i}+1}+(k-j)^{\alpha_{i}+1}-2(k-j+1)^{\alpha_{i}+1},&1\leq j\leq k,\\ 1,&j=k+1,\end{cases}$
and
$b^{i}_{j,k+1}=\frac{\Delta^{\alpha_{i}}}{\Gamma(\alpha_{i}+1)}\left((k+1-j)^{\alpha_{i}}-(k-j)^{\alpha_{i}}\right),\qquad 0\leq j\leq k.$
Here, $T$ denotes the terminal time or time horizon, $n$ the number of time steps, and $\Delta:=\frac{T}{N}$ the time increment.
Theoretical guarantees for the convergence of this numerical algorithm (the 2-dimensional fractional Adams–Bashforth–Moulton method) are established in Li and Tao (2009).
4.2 Numerical illustrations
Figure 1: Graph of the stabilizer $t\to\varsigma_{\alpha_{1},\lambda_{1},c_{1}}(t)$ (left) and 3030 samples paths $t_{k}\mapsto V^{1}_{t_{k}}$ (right) over the time interval $[0,1]$, for the Hurst esponent $H=0.4$, $c_{1}=0.01$ and number of time steps $n=600$.
Figure 2: Graph of $t_{k}\mapsto\text{Var}(V^{1}_{t_{k}},M)$ and $t_{k}\mapsto\mathbb{E}[\sigma^{2}(V^{1}_{t_{k}},M)]$ over $[0,1]$, $c_{1}=0.01$ and $n=600$.
Figure 3: Graph of $t_{k}\mapsto\psi^{1}_{t_{k}}$ and $t_{k}\mapsto\psi^{2}_{t_{k}}$ over $[0,1]$ with the fractional Adams algorithm, $\gamma=0.2$ and the number of time steps $n=200$ both for Power (left) and Exponential (right) utilities functions.
Since the optimal strategies $(\alpha^{*}_{t})_{t\in[0,T]}$ given by (3.51)–(3.79) are stochastic processes, we rather consider the evolution of the optimal vector of amounts invested in each stock, that is, the associated deterministic mapping $t\mapsto\pi_{t}^{*},$ (recall that $\alpha_{t}^{*}=\sigma(V_{t})^{\top}\pi_{t}^{*}$ with $\sigma(V_{t})=\sqrt{\mathrm{diag}(V_{t})}$, and $\alpha^{*}$ is given by (3.51)–(3.79)).
(A) $\gamma=0.2$
(B) $\gamma=0.5$
(C) $\gamma=0.8$
Figure 4: Evolution of the optimal portfolio strategy for different levels of the risk aversion parameter $\gamma$ for both the Power (left) and Exponential (right) utilities functions.
Plots show that, in addition to the impact of the roughness of asset volatility on the optimal allocation documented in Han and Wong (2020); Aichinger and Desmettre (2021), the stabilizing function also affects the optimal portfolio allocation, and hence the hedging demand over time.
5 Proofs of the main results
[…]
Lemma 5.1 (Martingale property of stochastic exponentials).
Let $W$, $W^{\top}$ be two independent $d$-dimensional Wiener processes and $V$ a weak solution to the inhomogeneous Volterra Square root Equation (2.2). Let $g_{1}$ and $g_{2}$ two deterministic processes with values in $\mathbb{R}^{d}$ such that $g_{1,i},g_{2,i}\in L^{\infty}(\mathbb{R}^{+},\mathbb{R})$ for $1\leq i\leq d$. Then the local martingale
$Z_{t}:=\mathcal{E}(\int_{0}^{t}\sum\limits_{i=1}^{d}g_{1,i}(s)\sqrt{V_{s}^{i}}dW^{i}_{s}+\int_{0}^{t}\sum\limits_{i=1}^{d}g_{2,i}(s)\sqrt{V_{s}^{i}}dW^{\top,i}_{s})$
is a true martingale.
Lemma 5.2 (Martingale property).
Let $g$ be a deterministic process with values in $\mathbb{R}$ and bounded such that $g\in L^{\infty}([0,T],\mathbb{R})$. Let us denote
${Z}_{t}:=\mathcal{E}(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s})=\exp\Big(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s}-\frac{1}{2}\int_{0}^{t}g^{2}(s)\left|\alpha_{s}\right|^{2}ds\Big),\quad t\in[0,T].$ (5.82)
where $(\alpha_{t})_{t\in[0,T]}$ is a $\mathbb{R}^{d}$- valued process belonging to $L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$}. Then $(Z_{t})_{t\in[0,T]}$ is a martingale.
Proof: This is a direct consequence of the fact that $g\in L^{\infty}([0,T],\mathbb{R})$ and $\alpha\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ together with Novikov’s condition.∎
We recall here the definition of one dimensional functional resolvent of the first kind of an integral kernel $K$ in the terminology of (B. Gripenberg and Saavalainen, 1990, Def. 5.5.1) (which can also be found e.g. in Abi Jaber et al. (2019)). Given $K\in{L}^{1}_{loc}(\mathbb{R}_{+})$, a function $r$ belonging to ${L}^{1}_{loc}(\mathbb{R}_{+})$ ( measure L on $\mathbb{R}_{+}$) is called functional resolvent of the first kind of $K$ if
$(K\star r)(t)=(r\star K)(t)=1,$ (5.83)
for all $t\in\mathbb{R}_{+}$. Note that the same notion can be defined for higher dimensions and in matrix form as follows: Let $K\in L^{1}_{\mathrm{loc}}(\mathbb{R}_{+},\mathbb{R}^{d\times d})$ and let $R$ be an $\mathbb{R}^{d\times d}$-valued measure on $\mathbb{R}_{+}$. Then $R$ is called the resolvent of the first kind of $K$ if
$K*R=R*K\equiv I,$ (5.84)
where $I$ denotes the $d$-dimensional identity matrix.
Lemma 5.3.
Let $f,g:\mathbb{R}_{+}\to\mathbb{R}$ be two locally bounded Borel function, let $K\!\in L^{1}_{loc}(Leb_{\mathbb{R}_{+}})$ and let $r\!\in L^{1}_{loc}(Leb_{\mathbb{R}_{+}})$ be its functional resolvent of the first kind. Then,
  1. $(a)$
    The Volterra Equation of the first kind $\forall\,t\geq 0,\quad f(t)=\int_{0}^{t}K(t-s)x(s)ds$ (also reading f=Kxf=K*x) has a solution given by:
    $\forall\,t\geq 0,\quad x(t)=\Big(\int_{0}^{t}f(t-s)r(ds)\Big)^{\prime}\;\quad\text{that is,}\quad x(t)=(f\star r)^{\prime}(t).$ (5.85)
    This solution is uniquely defined on $\mathbb{R}_{+}$ up to dtdt-$a.e.$ equality.
  2. $(b)$
    The resolvent kernel $r$ satisfies the following convolution identity:
    $\forall\,t\geq 0,\quad\int_{0}^{t}(f\star r)^{\prime}(t-s)(K\star g)(s)ds=\int_{0}^{t}f(t-s)g(s)ds=(f\star g)(t)$ (5.86)
We provide a proof of this classical result for the reader’s convenience.
Proof: The first claim in (5.85) follows by the associativity of the convolution and applying the fundamental theorem of calculus. For the second claim, we could use calculus of convolutions and resolvents. However, we rather used Laplace transform and deduce that:
$L_{\big((f\star r)^{\prime}\star(K\star g)\big)}(t)=tL_{f\star r}(t)L_{K\star g}(t)=tL_{f}(t)L_{r}(t)L_{K}(t)L_{g}(t)=L_{f}(t)L_{g}(t)=L_{f*g}(t).$
where the penultimalte equality come from applying laplace transform to equation 5.83. We next conclude by the injectivity of Laplace transform. $\Box$
5.1 Proofs of Proposition 3.4 and Theorems 3.5 and 3.7
Proof of Proposition 3.4 : We write for  $0\leq t\leq T$:
$\Gamma_{t}^{\frac{1}{\delta}}=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}\frac{\gamma}{\delta}\big(r(s)+\frac{\left|\lambda_{s}\right|^{2}}{2(1-\gamma)}\big)ds\Big)\mid\mathcal{F}_{t}\Big]=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}\frac{\gamma}{\delta}\big(r(s)+\frac{1}{2(1-\gamma)}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}\big)ds\Big)\mid\mathcal{F}_{t}\Big]$ (5.87)
which ensures that $\Gamma_{t}>0$ $\mathbb{P}-a.s.$, since $V_{t}$ is non-negative ($V\in\mathbb{R}^{d}_{+}$), $r(t)>0$ is deterministic, and $1-\gamma\leq\delta\leq 1$. We then have in view of (5.87) that there exists some positive constant $m>0$ such that $\Gamma_{t}\geq m>0$ for every $t\in[0,T]$. An application of the exponential-affine transform formula (Gnabeyeu, 2026, Theorem A.4.) with $\mathcal{M}\ni m(\,\mathrm{d}s):=\frac{\gamma}{2\delta(1-\gamma)}\theta\odot\theta\,{\rm Leb}_{d}(\,\mathrm{d}s)$ (where $\odot$ denote the Hadamard (pointwise or component-wise) product) yields:
$\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}\frac{\gamma}{2\delta(1-\gamma)}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big]=\exp\Big(\sum_{i=1}^{d}\int_{t}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2\delta(1-\gamma)}+F_{i}(s,\psi(T-s))\big)\tilde{g}^{i}_{t}(s)ds\Big)$
[…]
$\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\big(\int_{t}^{T}\frac{\gamma}{\delta}\big(r(s)+\frac{\left|\lambda_{s}\right|^{2}}{2(1-\gamma)}\big)ds\big)\mid\mathcal{F}_{t}\Big]=\exp\big(\frac{\gamma}{\delta}\int_{t}^{T}r(s)ds+\sum_{i=1}^{d}\int_{t}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2\delta(1-\gamma)}+F_{i}(s,\psi(T-s))\big)\tilde{g}^{i}_{t}(s)ds\big)$
This yields (3.36). Now, we set $G_{t}=\gamma\int_{t}^{T}r(s)ds+\sum_{i=1}^{d}\int_{t}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+\delta F_{i}(s,\psi(T-s))\big)\tilde{g}^{i}_{t}(s)ds,\quad t\leq T.$ Then, $\Gamma=\exp(G)$ and $d\Gamma_{t}=\Gamma_{t}\Big(dG_{t}+\frac{1}{2}d\langle G\rangle_{t}\Big).$ The dynamics of $G$ can readily be obtained by recalling $\tilde{g}_{t}(s)$ from (2.19) and by observing that for fixed $s$, the dynamics of $t\to\tilde{g}_{t}(s)$ are given by $d\tilde{g}_{t}(s)=K(s-t)\,d\tilde{Z}_{t}\quad t\leq s.$ Since $\tilde{g}_{t}(t)=V_{t}$, it follows by stochastic Fubini’s theorem, see Veraar (2012, Theorem 2.2), that the dynamics of $G$ reads as
$dG_{t}$ $=\;\Big(-\gamma r(t)-\sum_{i=1}^{d}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+\delta F_{i}(s,\psi(T-s))\big)V^{i}_{t}\Big)dt$
$\quad\quad+\delta\sum_{i=1}^{d}\int_{t}^{T}{K_{i}}(s-t)\big(\frac{\gamma\theta_{i}^{2}}{2\delta(1-\gamma)}+F_{i}(s,\psi(T-s))\big)dsd\tilde{Z}^{i}_{t}$
$=\;\Big(-\gamma r(t)-\sum_{i=1}^{d}\Big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+\delta\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(t)\psi^{i}(T-t))^{2}\Big)V^{i}_{t}\Big)dt+\;\delta\sum_{i=1}^{d}\psi^{i}(T-t)\nu_{i}\varsigma^{i}(t)\sqrt{V^{i}_{t}}d\widetilde{W}^{i}_{t},$
where we changed variables and used the inhomogeneous Riccati–Volterra equation (3.35) for $\psi$ for the last equality. This yields that the dynamics of $\Gamma$ is given by
$d\Gamma_{t}$ $=\;\Gamma_{t}\Big(-\Big(\gamma r(t)+\sum_{i=1}^{d}\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}V^{i}_{t}\Big)-\frac{\gamma}{1-\gamma}\delta^{2}\rho^{2}\sum_{i=1}^{d}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(t)\psi^{i}(T-t))^{2}V^{i}_{t}\Big)dt$
$\quad\quad+\;\delta\Gamma_{t}\sum_{i=1}^{d}\psi^{i}(T-t)\nu_{i}\varsigma^{i}(t)\sqrt{V^{i}_{t}}d\widetilde{W}^{i}_{t}$
$=\;\Gamma_{t}\Big[\big(-\gamma r(t)-\frac{\gamma}{2(1-\gamma)}\left|\lambda_{t}\right|^{2}-\frac{\gamma}{2(1-\gamma)}\delta^{2}\left|\Sigma\Lambda_{t}\right|^{2}\big)dt+\delta\Lambda_{t}^{\top}d\widetilde{W}_{t}\Big],$
where we used for the last identity the fact that
$\left|\Sigma\Lambda_{t}\right|^{2}=\;\sum_{i=1}^{d}\left(\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\right)^{2}V^{i}_{t}=\rho^{2}\sum_{i=1}^{d}\left(\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\right)^{2}V^{i}_{t},\quad\text{and}\quad\left|\lambda_{t}\right|^{2}=\;\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{t}$
Arguing as in the proof of Proposition 3.6, we show that $\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\sup_{t\in[0,T]}|\Gamma_{t}|^{p}\Big]<\infty$ for some $p>1$. As for $\Lambda$, it is clear that it belongs to $L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ for $\varsigma$ and $\psi$ are bounded and $\mathbb{E}\Big[\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\Big]<\infty$ thanks to (2.17). $\Box$
Proof of Theorem 3.5: We show that $J_{t}^{\alpha}$ fulfills the martingale optimality principle in Definition 3.2. For the first condition, note that $\Gamma_{T}=1$ and hence $J_{T}^{\alpha}=\frac{1}{\gamma}(X_{T}^{\alpha})^{\gamma}$. Since $\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$, $J^{\alpha}_{0}=\frac{x^{\gamma}_{0}}{\gamma}\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$ and consequently, the second condition is also satisfied. In order to show that the third condition is fulfilled, we apply Itô’s formula on $J_{t}^{\alpha}$. We have:
$dJ_{t}^{\alpha}=\;\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Gamma_{t}\Big[\big(-\gamma r(t)-\frac{\gamma}{2(1-\gamma)}(\left|\lambda_{t}\right|^{2}+\delta^{2}\left|\Sigma\Lambda_{t}\right|^{2})\big)dt+\delta\Lambda_{t}^{\top}d\widetilde{W}_{t}\Big]$
$\hskip 14.22636pt+\frac{(X_{t}^{\alpha})^{\gamma-1}}{\gamma}\Gamma_{t}\Big(\gamma X_{t}^{\alpha}\big(r(t)+\alpha_{t}^{\top}\lambda_{t}\big)+\frac{\gamma(\gamma-1)}{2}X_{t}^{\alpha}\alpha_{t}^{\top}\alpha_{t}\Big)dt+\Gamma_{t}(X_{t}^{\alpha})^{\gamma}\alpha_{t}^{\top}dB_{t}+\delta\alpha^{\top}_{t}\left(\Sigma\Lambda_{t}\right)(X_{t}^{\alpha})^{\gamma}\Gamma_{t}dt$
$\hskip 28.45274pt=\;J_{t}^{\alpha}\Big(\frac{\gamma(\gamma-1)}{2}\alpha^{\top}\alpha+\gamma\alpha_{t}^{\top}\lambda_{t}+\delta\alpha^{\top}_{t}\left(\Sigma\Lambda_{t}\right)-\frac{\gamma}{2(1-\gamma)}(\left|\lambda_{t}\right|^{2}+\delta^{2}\left|\Sigma\Lambda_{t}\right|^{2})-\delta\frac{\gamma}{(1-\gamma)}\Lambda_{t}^{\top}\Sigma\lambda_{t}\Big)dt$
$\hskip 28.45274pt+\gamma J_{t}^{\alpha}\alpha_{t}^{\top}dB_{t}+J_{t}^{\alpha}\delta\Lambda_{t}^{\top}dW_{t}=J_{t}^{\alpha}D_{t}(\alpha_{t})dt+\gamma J_{t}^{\alpha}\alpha_{t}^{\top}dB_{t}+J_{t}^{\alpha}\delta\Lambda_{t}^{\top}dW_{t}.$
where the drift factor takes the form:
$D_{t}(\alpha)=$ $\frac{\gamma(\gamma-1)}{2}\alpha^{\top}\alpha+\gamma\alpha^{\top}\big(\lambda_{t}+\delta\Sigma\Lambda_{t}\big)-\frac{\gamma}{2(1-\gamma)}\left|\lambda_{t}+\delta\Sigma\Lambda_{t}\right|^{2}.$
Differentiating $D_{t}(\alpha)$ with respect to $\alpha$ and checking the second order condition, one obtains the maximizer $\alpha^{*}_{t}=\frac{1}{1-\gamma}\left(\lambda_{t}+\delta\Sigma\Lambda_{t}\right)$ for every $t\in[0,T]$ that is the strategy given by Equation (3.39).
Evaluating the drift factor $D_{t}$ at $\alpha^{*}_{t}$ show that $D_{t}(\alpha^{*}_{t})$ vanishes to 0. Note $D_{t}(\alpha)$ is a quadratic function on $\alpha$ and $\gamma-1<0$. As $D_{t}(\alpha^{*}_{t})=0$, then $D_{t}(\alpha_{t})\leq 0$ for any admissible strategy $\alpha$. Moreover, solving the stochastic differential equation for $J_{t}^{\alpha}$ yields
$\hskip 142.26378pt\forall\,t\in[0,T]\quad J_{t}^{\alpha}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}e^{\int_{0}^{t}D_{s}(\alpha_{s})\,ds}F_{t}^{\alpha}\quad\text{where}$ (5.88)
$F_{t}^{\alpha}=\mathcal{E}\big(\int_{0}^{t}\gamma\,\alpha_{s}^{\top}\,dB_{s}+\delta\,\Lambda_{s}^{\top}\,dW_{s}\big)=\operatorname{exp}\Big(-\frac{1}{2}\int^{t}_{0}\big(\gamma^{2}\left|\alpha_{s}\right|^{2}+\delta^{2}\left|\Lambda_{s}\right|^{2}\big)ds+\gamma\int_{0}^{t}\alpha_{s}^{\top}dB_{s}+\delta\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big).$
Now, since $D_{s}(\alpha_{s})\leq 0$, $e^{\int^{t}_{0}D_{s}(\alpha_{s})ds}$ is a non-increasing function. By our assumptions on the admissible strategies 3.3 and Proposition 3.4, $\left(\alpha,\Lambda\right)\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ and thus the stochastic exponential $F_{t}^{\alpha}$ is a local martingale (which follows from the basic properties of the Dooléans-Dade exponential). Therefore, there exists a sequence of stopping times $\{\tau_{n}\}_{n\geq 1}$ satisfying $\lim_{n\rightarrow\infty}\tau_{n}=T$, $\mathbb{P}$-a.s., such that
$\mathbb{E}[J^{\alpha}_{t\wedge\tau_{n}}|{\mathcal{F}}_{s}]\leq J^{\alpha}_{s\wedge\tau_{n}},\quad s\leq t\leq T,$
for every $n$. Moreover, since $J^{\alpha}_{t}$ is bounded from below by 0 ( $J_{t}^{\alpha}\geq 0$), applying Fatou’s Lemma for $n\rightarrow\infty$, we deduce that $J^{\alpha}_{t}$ is a supermartingale for every arbitrary admissible strategy $\alpha$. It remains to show that $J_{t}^{\alpha^{*}}$ is a true martingale for the optimal strategy $\alpha^{*}$ in which case $e^{\int_{0}^{t}D_{s}(\alpha_{s}^{*})ds}=1$ and hence $J_{t}^{\alpha^{*}}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}F_{t}^{\alpha^{*}}$.
For $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a martingale by Lemma 5.1 with $g_{1,i}(t)=\frac{\gamma}{1-\gamma}\rho_{i}\theta_{i}+(1+\frac{\gamma}{1-\gamma}\rho_{i}^{2})\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\in L^{\infty}([0,T],\mathbb{R})$ and $g_{2,i}(t)=\frac{\gamma}{1-\gamma}\sqrt{1-\rho_{i}^{2}}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\in L^{\infty}([0,T],\mathbb{R})$ for $1\leq i\leq d$. Subsequently, $J^{\alpha^{*}}_{t}$ is a true martingale. We have verified all conditions required by martingale optimality principle, except for the admissibility of $\alpha^{*}$, which follows directly from Step 1 in the proof of Theorem 3.7 in the more general setting, as the arguments are similar. The proof is complete. $\Box$
Proof of Theorem 3.7:
Step 1 (Proof of Admissibility:) We begin by proving the admissibility of the candidate in equation (3.49) for the optimal portfolio strategy $\alpha^{*}$, and, en route, we deduce the bound in equation (3.52).
In order to show that $\alpha^{*}$ is admissible, we have to show that (a), (b), (c) of Definition 3.3 hold. Part (a) is true because (3.30) has a unique solution (3.31) in terms of $(S,V,B)$. For part (b) it suffices to show that $\mathbb{E}[\operatorname{sup}_{t\in[0,T]}|X_{t}^{\alpha^{*}}|^{p}]<\infty$ for every $p>0$. Inserting the explicit solution of the control problem into (3.30) and using Itô’s lemma (or equivalently into (3.31)), the unique continuous solution of the controlled wealth process $X^{\alpha^{*}}$ is given by :
$X_{t}^{\alpha^{*}}=x_{0}\exp\!\left(\int_{0}^{t}\Big(r(s)+\alpha_{s}^{*\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}^{*}\right|^{2}\Big)\,ds+\int_{0}^{t}\alpha_{s}^{*\top}\,dB_{s}\right).$ (5.89)
Observe by virtue of (3.27), that the Doléans-Dade exponential $\mathcal{E}\left(\int_{0}^{\cdot}\alpha_{s}^{*\top}dB_{s}\right)$ satisfies Novikov’s condition, and is therefore a true martingale (see also Lemma 5.2). By virtue of the Cauchy-Schwarz inequality, and then Doob’s maximal inequality in the second line, for every $p>1$
$\,\mathbb{E}\Big[\sup_{t\in[0,T]}|X^{\alpha^{*}}_{t}|^{p}\Big]\leq x_{0}^{p}\mathbb{E}\Big[\sup_{t\in[0,T]}\big|e^{\int_{0}^{t}{\left(r(s)+\lambda_{s}^{\top}\alpha_{s}^{*}\right)}ds}\big|^{2p}\Big]^{\frac{1}{2}}\mathbb{E}\Big[\sup_{t\in[0,T]}\Big|\exp\Big(-\tfrac{1}{2}\int^{t}_{0}\left|\alpha_{s}^{*}\right|^{2}ds+\int^{t}_{0}\alpha_{s}^{*\top}\,dB_{s}\Big)\Big|^{2p}\Big]^{\frac{1}{2}}$
$\hskip 65.44142pt\leq x_{0}^{p}e^{p\int_{0}^{T}r(s)ds}\Big(\frac{2p}{2p-1}\Big)^{p}\mathbb{E}\Big[e^{2p\int_{0}^{T}|\lambda_{s}^{\top}\alpha_{s}^{*}|ds}\Big]^{\frac{1}{2}}\mathbb{E}\Big[\exp\Big(-p\int^{T}_{0}\left|\alpha_{s}^{*}\right|^{2}ds+2p\int^{T}_{0}\alpha_{s}^{*\top}\,dB_{s}\Big)\Big]^{\frac{1}{2}}.$
The first term is finite on the first hand by condition (3.27) with constant $a(\frac{p}{1-\gamma})=\frac{p}{1-\gamma}\left(2+|\Sigma|\right)$ and owing to the elementary arithmetic mean-geometric mean (AM-GM) inequality $|ab|\leq(|a|^{2}+|b|^{2})/2$
$\mathbb{E}\Big[e^{2p\int_{0}^{T}|\lambda_{s}^{\top}\alpha_{s}^{*}|ds}\Big]\leq\mathbb{E}\left[\exp\left({a(\frac{p}{1-\gamma})\int_{0}^{T}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds}\right)\right]<\infty,$ (5.90)
and, on the other hand, the second term is also finite since by virtue of Hölder’s inequality,
$\mathbb{E}\Big[e^{-\int^{T}_{0}p\left|\alpha_{s}^{*}\right|^{2}ds+\int^{T}_{0}2p\alpha_{s}^{*\top}\,dB_{s}}\Big]$ $\leq\left(\mathbb{E}\Big[e^{(8p^{2}-2p)\int^{T}_{0}\left|\alpha_{s}^{*}\right|^{2}ds}\Big]\right)^{1/2}\left(\mathbb{E}\Big[e^{-8p^{2}\int^{T}_{0}\left|\alpha_{s}^{*}\right|^{2}ds+4p\int^{T}_{0}\alpha_{s}^{*\top}\,dB_{s}}\Big]\right)^{1/2}$
$\leq\left(\mathbb{E}\left[e^{a(\frac{p}{1-\gamma})\int_{0}^{T}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds}\right]\right)^{1/2}\times 1<\infty.$
with constant $a(\frac{p}{1-\gamma})=2(8\big(\frac{1}{1-\gamma}\big)^{2}{-2\frac{p}{1-\gamma}})\left(1+|{\Sigma}|^{2}\right)$ and where we used Jensen’s inequality to bound
$\left|\alpha_{s}^{*}\right|^{2}=\big(\frac{1}{1-\gamma}\big)^{2}|\lambda_{s}+\Sigma\Lambda_{s}|^{2}\leq 2\big(\frac{1}{1-\gamma}\big)^{2}(|\lambda_{s}|^{2}+|\Sigma\Lambda_{s}|^{2})\leq 2\big(\frac{1}{1-\gamma}\big)^{2}(1+|\Sigma|^{2})(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}).$ (5.91)
then noticing that $\big(\frac{1}{1-\gamma}\big)^{2}(8p^{2}{-2p})\leq 8\big(\frac{p}{1-\gamma}\big)^{2}{-2\frac{p}{1-\gamma}}$ for all $p>1,\,\gamma\in(0,1)$, together with condition (3.27) and Novikov’s condition to the Doléans-Dade exponential $\mathcal{E}(4p\int_{0}^{{\cdot}}\alpha_{s}^{*\top}dB_{s})$. This leads to Part (c). $\mathbb{E}\Big[\sup_{t\in[0,T]}|X^{\alpha^{*}}_{t}|^{p}\Big]<\infty$ is proved. It becomes straightforward to verify $\alpha^{*}$ is admissible. We are left to prove that $\alpha^{*}\in L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. We have
$\mathbb{E}\left[\int_{0}^{T}|\alpha_{s}^{*}|^{2}ds\right]$ ${=}\mathbb{E}\left[\int_{0}^{T}\big(\frac{1}{1-\gamma}\big)^{2}|\lambda_{s}+\Sigma\Lambda_{s}|^{2}ds\right]\leq\mathbb{E}\left[2\big(\frac{1}{1-\gamma}\big)^{2}(1+|\Sigma|^{2})\int_{0}^{T}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds\right]<\infty,$
where the last term is finite due to condition (3.27) and the inequality $|z|^{q}\leq c_{q}e^{|z|},\;\forall\,q\geq 1$.
Step 2 (Proof of the equality (1):) Let’s us consider the Problem (3.33) with an arbitrary strategy $\alpha\in\mathcal{A}$ from Definition 3.3. To ease notations, we set $h_{t}=\lambda_{t}+\Sigma\Lambda_{t}$. For any admissible strategy $\alpha\in\mathcal{A}$, using the Riccati BSDE (3.46) in Proposition 3.6 together with Itô’s lemma yield:
$d\big(\frac{(X^{\alpha}_{t})^{\gamma}}{\gamma}\Gamma_{t}\big)\;$ $=\frac{(X^{\alpha}_{t})^{\gamma}}{\gamma}\Gamma_{t}\Big(D_{t}(\alpha_{t})dt+\gamma\alpha_{t}^{\top}dB_{t}+\Lambda_{t}^{\top}dW_{t}\Big).$
where the drift factor takes the form $D_{t}(\alpha_{t})=\frac{\gamma(\gamma-1)}{2}\alpha_{t}^{\top}\alpha_{t}+\gamma\alpha_{t}^{\top}h_{t}-\frac{\gamma}{2(1-\gamma)}h_{t}^{\top}h_{t}.$ As a consequence, using $\Gamma_{T}=1$, we get
$\frac{(X^{\alpha}_{T})^{\gamma}}{\gamma}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}e^{\int^{T}_{0}D_{s}(\alpha_{s})ds}F_{T}^{\alpha}.$ (5.92)
where $F_{T}^{\alpha}$ is an $(\mathbb{F},\mathbb{P})$-local martingale as $(\alpha,\Lambda)$ are in $L^{2,{loc}}_{\mathbb{F}}([0,T])$ (Dooléans-Dade exponential) and write
$F_{t}^{\alpha}=\mathcal{E}\Big(\int_{0}^{t}\gamma\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)=\operatorname{exp}\Big(-\frac{1}{2}\int^{t}_{0}\Big(\gamma^{2}\left|\alpha_{s}\right|^{2}+\left|\Lambda_{s}\right|^{2}\Big)ds+\gamma\int_{0}^{t}\alpha_{s}^{\top}dB_{s}+\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big).$
At this stage, for $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a $\mathbb{P}$-martingale with expectation $1$ by Lemma 5.1 with $g_{1,i}(t)=\frac{\gamma}{1-\gamma}\rho_{i}\theta_{i}+(1+\frac{\gamma}{1-\gamma}\rho_{i}^{2})\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\in L^{\infty}([0,T],\mathbb{R})$ and $g_{2,i}(t)=\frac{\gamma}{1-\gamma}\sqrt{1-\rho_{i}^{2}}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\in L^{\infty}([0,T],\mathbb{R})$ for $1\leq i\leq d$. Consequently, inserting the candidate (3.49) for the optimal strategy $\alpha^{*}$ in the above Equation (5.92), and observing that $D_{s}(\alpha_{s}^{*})=0\;\forall\,s\in[0,T]$, then taking the expectation, it is straightforward that:
$\mathbb{E}_{x_{0},V_{0}}{\big[\frac{(X^{\alpha^{*}}_{T})^{\gamma}}{\gamma}\big]}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}\mathbb{E}_{x_{0},V_{0}}{\big[F_{T}^{\alpha^{*}}\big]}=\frac{x_{0}^{\gamma}}{\gamma}\Gamma_{0}.$
where the last inequality comes from the martingality of $F_{T}^{\alpha^{*}}$ and the desired result about the first part of the proof (equality (1)) is completed. It remains to show the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$.
Step 3 (Proof of the inequality (2):) Recall that the SDE for the wealth process (3.30) can be solved explicitly so as to admit the representation
$X_{T}^{\alpha}=X_{0}^{\alpha}\exp\!\left(\int_{0}^{T}\Big(r(s)+\alpha_{s}^{\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}\right|^{2}\Big)\,ds+\int_{0}^{T}\alpha_{s}^{\top}\,dB_{s}\right)\quad\text{with}\quad X_{0}^{\alpha}=x_{0}.$ (5.93)
Since $\alpha_{t}$ is bounded by assumption, we can introduce a new probability measure $\mathbb{Q}$ with Radon-Nikodym density or derivative at $F_{T}$ as
${Z}_{T}:=\frac{d{\mathbb{Q}}}{d\mathbb{P}}|_{\mathcal{F}_{T}}=\exp\Big(\gamma\int_{0}^{T}\alpha_{s}^{\top}\,dB_{s}-\frac{\gamma^{2}}{2}\int_{0}^{T}\left|\alpha_{s}\right|^{2}ds\Big),$ (5.94)
by Lemma 5.2. By Girsanov’s theorem, we may write:
$\,(x_{0})^{-\gamma}\mathbb{E}_{x_{0},V_{0}}\Big[(X_{T}^{\alpha})^{\gamma}\Big]=\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}\gamma r(s)ds+\gamma\int_{0}^{T}\big(\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma-1}{2}\left|\alpha_{s}\right|^{2}\big)ds\Big)\Big],$
$\hskip 102.43008pt=\exp\Big(\gamma\int_{0}^{T}r(s)ds\Big)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}ds\Big)\Big],$
with the real-valued deterministic process $F_{s}$ given by
$\forall\,s\in[0,T],\quad F_{s}^{\alpha}=\gamma\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma(\gamma-1)}{2}\left|\alpha_{s}\right|^{2}=F_{s}^{\alpha^{*}}+\gamma\hat{\alpha}_{s}^{\top}\lambda_{s}+\gamma(\gamma-1)\alpha_{s}^{*\top}\hat{\alpha}_{s}+\frac{\gamma(\gamma-1)}{2}\left|\hat{\alpha}_{s}\right|^{2}$
where we write the arbitrary admissible strategy $(\alpha_{t})_{t\in[0,T]}$ in terms of the optimal strategy $\alpha^{*}_{t}=\frac{1}{1-\gamma}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)$ for every $t\in[0,T]$ and some remainder $(\hat{\alpha}_{t})_{t\in[0,T]}$ i.e. $\alpha_{t}=\alpha_{t}^{*}+\hat{\alpha}_{t}$ for every $t\in[0,T]$. Note that, on the first hand, substituting the optimal strategy $\alpha^{*}$ in (3.49) yields $\forall\,s\in[0,T]$
$F_{s}^{\alpha^{*}}=\gamma\big(\alpha_{s}^{*\top}\lambda_{s}+\frac{\gamma-1}{2}\left|\alpha_{s}^{*}\right|^{2}\big)=\frac{\gamma}{2(1-\gamma)}\big(\left|\lambda_{s}\right|^{2}-\left|\Sigma\Lambda_{s}\right|^{2}\big)=\frac{\gamma}{2(1-\gamma)}\sum_{i=1}^{d}\Big(\theta_{i}^{2}-\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}$
and on the other hand, recalling that $\alpha_{s}:=\sqrt{\operatorname{diag}(V_{s})}^{\top}\pi_{s}$ and defining $\hat{\pi}:=\pi-\pi^{*}$ so that $\pi_{s}=\pi_{s}^{*}+\hat{\pi}_{s}$, with $\pi_{s}^{*}=\frac{1}{1-\gamma}\left(\theta+\Sigma\nu\varsigma(s)\psi(T-s)\right)$ so that $\pi_{s,i}^{*}=\frac{1}{1-\gamma}(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s))$ for $i=1,\ldots,d$ and consequently, $\forall\,s\in[0,T],$
$F_{s}^{\alpha}$ $=F_{s}^{\alpha^{*}}+\sum_{i=1}^{d}\Big(\gamma\theta_{i}\hat{\pi}_{s,i}+\gamma(\gamma-1)\pi_{s,i}^{*}\hat{\pi}_{s,i}+\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
$=\sum_{i=1}^{d}\Big(\frac{\gamma}{2(1-\gamma)}\Big(\theta_{i}^{2}-\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)-\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}+\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
Using Equation (3.43) reading $\frac{\gamma}{2(1-\gamma)}\theta_{i}^{2}+F_{i}(s,\psi(T-s))=(\psi^{i}\star r^{i})^{\prime}(T-s)$ owing to Lemma 5.3 $(a)$ for $i=1,\ldots,d$, where $r^{i}$ the resolvent of the first kind of the integral kernel $K_{i}$, one gets:
$\forall\,s\in[0,T],\;F_{s}$ $=\sum_{i=1}^{d}(\psi^{i}\star r^{i})^{\prime}(T-s)V^{i}_{s}-\Big(F_{i}(s,\psi(T-s))+\frac{\gamma\rho_{i}^{2}\nu_{i}^{2}}{2(1-\gamma)}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}$ (5.95)
$\hskip 42.67912pt+\sum_{i=1}^{d}\Big(-\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}+\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
Under the probability measure $\mathbb{Q}$ defined in (5.105), the processes $\widehat{B}_{t}:=B_{t}-\gamma\int_{0}^{t}\alpha_{s}ds$ and
$\widehat{W}_{t}:=\Sigma\widehat{B}_{t}+\sqrt{I-\Sigma^{\top}\Sigma}B^{\perp}_{t}=W_{t}-\gamma\int_{0}^{t}\Sigma\alpha_{s}ds=W_{t}-\gamma\int_{0}^{t}\Sigma\alpha^{*}_{s}ds-\gamma\int_{0}^{t}\Sigma\hat{\alpha}_{s}ds,$
are standard Wiener processes thanks to the Cameron–Martin–Girsanov and Girsanov theorem respectively theorem. The dynamics of the variance process $V$ under $\mathbb{Q}$ can thus be written as
$\,V_{t}^{i}=g_{0}^{i}(t)+\int_{0}^{t}K_{i}(t-s)H_{s}^{i}ds+\int_{0}^{t}K_{i}(t-s)\nu_{i}\varsigma^{i}(s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s},\quad i=1,\ldots,d;\quad\text{with}$
$H_{s}^{i}:=(DV_{s})_{i}+\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)(\pi_{s,i}^{*}+\hat{\pi}_{s,i})V_{s}^{i},\quad i=1,\ldots,d.$
$\hskip 17.07182pt=\frac{\gamma}{1-\gamma}\theta_{i}\rho_{i}\nu_{i}\varsigma^{i}(s)V_{s}^{i}+(DV_{s})_{i}+\frac{\gamma}{1-\gamma}(\rho_{i}\nu_{i}\varsigma^{i}(s))^{2}\psi^{i}(T-s)V_{s}^{i}+\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)\hat{\pi}_{s,i}V_{s}^{i}$
Now, we insert the dynamics of $V^{i}$ into the expression $\int_{t}^{T}(\psi^{i}\star r^{i})^{\prime}(T-s)V^{i}_{s}ds$ and simplify using Lemma 5.3 $(b)$, where we reads $H^{i}$ and the stochastic term pathwise (i.e. $\omega$ by $\omega$) as for example $g(s)=H^{i}_{s}(\omega)$. Its boils down that:
$\,\int_{t}^{T}\sum_{i=1}^{d}(\psi^{i}\star r^{i})^{\prime}(T-s)V^{i}_{s}ds=\sum_{i=1}^{d}\int_{t}^{T}(\psi^{i}\star r^{i})^{\prime}(T-s)g_{0}^{i}(s)ds+\int_{t}^{T}\psi^{i}(T-s)\big(H_{s}^{i}ds+\nu_{i}\varsigma^{i}(s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\big)$
$\hskip 42.67912pt=\sum_{i=1}^{d}\int_{t}^{T}\big(\frac{\gamma}{2(1-\gamma)}\theta_{i}^{2}+F_{i}(s,\psi(T-s))\big)g_{0}^{i}(s)ds+\sum_{i=1}^{d}\int_{t}^{T}\psi^{i}(T-s)\big(H_{s}^{i}ds+\nu_{i}\varsigma^{i}(s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\big).$
Noting by a change of variables that
$\sum_{j=1}^{d}\psi^{j}(T-s)(DV_{s})_{j}=\sum_{j=1}^{d}\psi^{j}(T-s)\sum_{i=1}^{d}D_{ji}V_{s}^{i}=\sum_{i=1}^{d}V_{s}^{i}\sum_{j=1}^{d}D_{ji}\psi^{j}(T-s)=\sum_{i=1}^{d}(D^{\top}\psi)_{i}(T-s)V_{s}^{i},$
one gets using Equation (3.44)
$\sum_{i=1}^{d}\int_{0}^{T}\psi^{i}(T-s)H_{s}^{i}ds$ $=\sum_{i=1}^{d}\int_{0}^{T}\Big(F_{i}(s,\psi^{i}(T-s))+\big(\frac{\gamma}{2(1-\gamma)}\rho_{i}^{2}-\frac{1}{2}\big)\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V_{s}^{i}ds$
$+\gamma\sum_{i=1}^{d}\int_{0}^{T}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}V_{s}^{i}ds.$
Replacing back all in (5.95), some terms cancel out and we end up with
$\int_{0}^{T}F_{s}\,ds$ $=\sum_{i=1}^{d}\int_{0}^{T}\Big(\frac{\gamma}{2(1-\gamma)}\theta_{i}^{2}+F_{i}(s,\psi(T-s))\Big)g_{0}^{i}(s)\,ds+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds$
$\qquad+\sum_{i=1}^{d}\int_{0}^{T}\Big(-\frac{\nu_{i}^{2}}{2}\big(\varsigma^{i}(s)\psi^{i}(T-s)\big)^{2}V_{s}^{i}\,ds+\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}\,d\widehat{W}^{i}_{s}\Big).$
Hence we obtain
$\,x_{0}^{-\gamma}\mathbb{E}_{x_{0},V_{0}}[{(X_{T}^{\alpha})^{\gamma}}]=\exp\Big(\gamma\int_{0}^{T}r(s)ds\Big)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}^{\alpha}ds\Big)\Big],$
$\hskip 99.58464pt=\exp\Big(\gamma\int_{0}^{T}r(s)ds+\sum_{i=1}^{d}\int_{0}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 79.6678pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}$
$\hskip 170.71652pt+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds\Big)\Big].$ (5.96)
Remark: (Alternative Proof of the equality (1):) At this stage, note that if the admissible strategy $\alpha$ is optimal, i.e. $\alpha=\alpha^{*}$ so that $\hat{\pi}\equiv 0$ and $\mathbb{Q}\equiv\mathbb{Q}^{*}$ in (5.94), then Equation (5.96) becomes:
$x_{0}^{-\gamma}\mathbb{E}_{x_{0},V_{0}}[{(X_{T}^{\alpha^{*}})^{\gamma}}]=\exp\Big(\gamma\int_{0}^{T}r(s)ds+\sum_{i=1}^{d}\int_{0}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 59.75095pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}^{*}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{*i}_{s}\Big)\Big].$
Since $\psi$ is continuous, it is bounded and therefore the stochastic exponential is a true $\mathbb{Q}^{*}$- martingale with expectation $1$ by Lemma 5.1 with $g_{2}=0$ and $g_{1,i}(s)=\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\in L^{\infty}(\mathbb{R}^{+},\mathbb{R})$ . Thus we get
$\mathbb{E}_{x_{0},V_{0}}\Big[{\frac{1}{\gamma}(X_{T}^{\alpha^{*}})^{\gamma}}\Big]=\frac{x_{0}^{\gamma}}{\gamma}\exp\Big(\gamma\int_{0}^{T}r(s)ds+\sum_{i=1}^{d}\int_{0}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$
This again completes the first part of the proof (equality (1)).
Coming back to the proof of the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$, from Equation (5.96), since $V_{s}$ is positive definite and $\gamma\in(0,1)$, the term $\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds=\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma(\gamma-1)}{2}\hat{\alpha}_{s,i}^{2}\,ds$ has to be less than or equal to 0 so that
$x_{0}^{-\gamma}\mathbb{E}_{x_{0},V_{0}}[{(X_{T}^{\alpha})^{\gamma}}]\leq\exp\Big(\gamma\int_{0}^{T}r(s)ds+\sum_{i=1}^{d}\int_{0}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 62.59596pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\Big)\Big].$
Since the stochastic exponential is a $\mathbb{Q}$-martingale with expectation $1$ still by Lemma 5.1, we finally obtain that for every $\alpha\in\mathcal{A}$
$\mathbb{E}_{x_{0},V_{0}}\Big[{\frac{1}{\gamma}(X_{T}^{\alpha})^{\gamma}}\Big]\leq\frac{x_{0}^{\gamma}}{\gamma}\exp\Big(\gamma\int_{0}^{T}r(s)ds+\sum_{i=1}^{d}\int_{0}^{T}\big(\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$
which completes the proof. $\Box$
5.2 Proofs of Proposition 3.9 and Theorems 3.10 and 3.12
[…]
$\Gamma_{t}^{1-\rho^{2}}=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\Big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\left|\lambda_{s}\right|^{2}ds\Big)|\mathcal{F}_{t}\Big]=\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big],$ (5.97)
which ensures that $\Gamma_{t}\leq 1$, $\tilde{\mathbb{P}}-a.s.$, since $V\in\mathbb{R}^{d}_{+}$. An application of the exponential-affine transform formula in (Gnabeyeu, 2026, Theorem A.4) with $\mathcal{M}\ni m(\,\mathrm{d}s):=-\frac{1-\rho^{2}}{2}\theta\odot\theta\,{\rm Leb}_{d}(\,\mathrm{d}s)$ (where $\odot$ denote the Hadamard (pointwise or component-wise) product) yields:
$\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}-\frac{1-\rho^{2}}{2}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big]=\exp\Big(\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{1-\rho^{2}}{2}\theta_{i}^{2}+\tilde{F}_{i}(s,\tilde{\psi}(T-s))\big)\tilde{g}^{i}_{t}(s)ds\Big)$ (5.98)
[…]
$\tilde{F}_{i}(s,\tilde{\psi})=-\rho\theta_{i}\nu_{i}\varsigma^{i}(s)\tilde{\psi}^{i}+(D^{\top}\tilde{\psi})_{i}+\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\tilde{\psi}^{i})^{2},\quad i=1,\ldots,d,$
and $\tilde{\psi}\in C([0,T],(\mathbb{R}^{d})^{*})$ solves the inhomogeneous Ricatti-Volterra equation
$\tilde{\psi}^{i}(t)$ $=\int_{0}^{t}K_{i}(t-s)\big(-\frac{\theta_{i}^{2}(1-\rho^{2})}{2}+\tilde{F}_{i}(T-s,\tilde{\psi}(s))\big)ds,\quad i=1,\ldots,d.$ (5.99)
Setting $\tilde{\psi}=(1-\rho^{2})\psi$ implies that $\tilde{F}_{i}(s,\tilde{\psi})=(1-\rho^{2})F_{i}(s,\psi)\quad i=1,\ldots,d$, where $F_{i}$ is given in (3.58). Therefore, it holds that for all $t\in[0,T]$,
$\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\exp\Big(\int_{t}^{T}-\frac{1-\rho^{2}}{2}\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{s}ds\Big)\mid\mathcal{F}_{t}\Big]=\exp\Big((1-\rho^{2})\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)\tilde{g}^{i}_{t}(s)ds\Big)$
Consequently, (3.59) holds and $\psi\in L^{2}([0,T],(\mathbb{R}^{d})^{*})$ solves the inhomogeneous Ricatti-Volterra equation (3.57)- (3.58). Note that for every $t\in[0,T]$, Equation (3.59) implies straightforwardly that $\Gamma_{t}>0$, $\mathbb{P}-a.s.$. Now, we set
$G_{t}=\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)\tilde{g}^{i}_{t}(s)ds,\quad t\leq T.$
Then, $\Gamma=\exp(G)$ and $d\Gamma_{t}=\Gamma_{t}\Big(dG_{t}+\frac{1}{2}d\langle G\rangle_{t}\Big).$ The dynamics of $G$ can readily be obtained by recalling $\tilde{g}_{t}(s)$ from (2.19) and by observing that for fixed $s$, the dynamics of $t\to\tilde{g}_{t}(s)$ are given by $d\tilde{g}_{t}(s)=K(s-t)\,d\tilde{Z}_{t}\quad t\leq s.$ Since $\tilde{g}_{t}(t)=V_{t}$, it follows by stochastic Fubini’s theorem, see Veraar (2012, Theorem 2.2), that the dynamics of $G$ reads as
$dG_{t}$ $=\;\Big(\sum_{i=1}^{d}\big(\frac{\theta_{i}^{2}}{2}-F_{i}(s,\psi(T-s))\big)V^{i}_{t}\Big)dt+\sum_{i=1}^{d}\int_{t}^{T}{K_{i}}(s-t)\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)dsd\tilde{Z}^{i}_{t}$
$=\;\Big(\sum_{i=1}^{d}\Big(\frac{\theta_{i}^{2}}{2}-\frac{\nu_{i}^{2}}{2}(1-\rho^{2})(\varsigma^{i}(t)\psi^{i}(T-t))^{2}\Big)V^{i}_{t}\Big)dt+\sum_{i=1}^{d}\psi^{i}(T-t)\nu_{i}\varsigma^{i}(t)\sqrt{V^{i}_{t}}d\widetilde{W}^{i}_{t},$
where we changed variables and used the Riccati–Volterra equation (3.58) for $\psi$ for the last equality. This yields that the dynamics of $\Gamma$ is given by
$d\Gamma_{t}$ $=\;\Gamma_{t}\Big(\sum_{i=1}^{d}\Big(\frac{\theta_{i}^{2}}{2}-\frac{\nu_{i}^{2}}{2}(1-\rho^{2})(\varsigma^{i}(t)\psi^{i}(T-t))^{2}\Big)V^{i}_{t}+\sum_{i=1}^{d}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(t)\psi^{i}(T-t))^{2}V^{i}_{t}\Big)dt$
$\quad\quad+\;\Gamma_{t}\sum_{i=1}^{d}\psi^{i}(T-t)\nu_{i}\varsigma^{i}(t)\sqrt{V^{i}_{t}}d\widetilde{W}^{i}_{t}=\;\Gamma_{t}\Big[\big(\frac{1}{2}\left|\lambda_{t}\right|^{2}+\frac{1}{2}\left|\Sigma\Lambda_{t}\right|^{2}\big)dt+\Lambda_{t}^{\top}d\widetilde{W}_{t}\Big],$
where we used for the last identity the fact that $\left|\Sigma\Lambda_{t}\right|^{2}=\;\rho^{2}\sum_{i=1}^{d}\left(\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\right)^{2}V^{i}_{t},\;\text{and}\;\left|\lambda_{t}\right|^{2}=\;\sum_{i=1}^{d}\theta_{i}^{2}V^{i}_{t}.$ As $0<\Gamma_{t}\leq 1$, $\mathbb{P}-a.s.$, we have that $\Gamma\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})$. As for $\Lambda$, it is clear that it belongs to $L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ since $\varsigma$ and $\psi$ are bounded, $\psi$ is continuous thus bounded and $\mathbb{E}\Big[\int_{0}^{T}\sum_{i=1}^{d}V^{i}_{s}ds\Big]<\infty$ by (2.17). Consequently, $\left(\Gamma,\Lambda\right)\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. This complete the Proof $\Box$
Proof of Theorem 3.10 : We show that $J_{t}^{\alpha}$ fulfills the martingale optimality principle. For the first condition, note that $\Gamma_{T}=1$ and hence $J_{T}^{\alpha}=-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)$. Since $\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$, $J^{\alpha}_{0}=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$ and thus the second condition is also satisfied. In order to show that the third condition is also fulfilled, we apply Itô’s formula on $J_{t}^{\alpha}$. Setting $Y_{t}:=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\right)$, we write by Itô’s lemma:
$dY_{t}$ $=\left(\gamma r(s)e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\big(r(t)X^{\alpha}_{t}+\alpha_{t}^{\top}\lambda_{t}\big)+\frac{\gamma^{2}}{2}\alpha_{t}^{\top}\alpha_{t}e^{\int_{t}^{T}2r(s)ds}\right)Y_{t}dt-\gamma e^{\int_{t}^{T}r(s)ds}Y_{t}\alpha_{t}^{\top}dB_{t}$
$=\left(-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}\lambda_{t}+\frac{\gamma^{2}}{2}\alpha_{t}^{\top}\alpha_{t}e^{\int_{t}^{T}2r(s)ds}\right)Y_{t}dt-\gamma e^{\int_{t}^{T}r(s)ds}Y_{t}\alpha_{t}^{\top}dB_{t}$
Consequently, by Itô’s product rule, one may write
$dJ_{t}^{\alpha}=$ $\;Y_{t}\Gamma_{t}\Big[\big(\frac{1}{2}\left|\lambda_{t}\right|^{2}+\frac{1}{2}\left|\Sigma\Lambda_{t}\right|^{2}\big)dt+\Lambda_{t}^{\top}d\widetilde{W}_{t}\Big]+Y_{t}\Gamma_{t}\left(-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}\lambda_{t}+\frac{\gamma^{2}}{2}\alpha_{t}^{\top}\alpha_{t}e^{\int_{t}^{T}2r(s)ds}\right)dt$
$-Y_{t}\Gamma_{t}\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\alpha^{\top}_{t}\left(\Sigma\Lambda_{t}\right)Y_{t}\Gamma_{t}dt$
$=\;J_{t}^{\alpha}\Big(\frac{\gamma^{2}}{2}e^{\int_{t}^{T}2r(s)ds}\alpha^{\top}_{t}\alpha_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}\lambda_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\alpha^{\top}_{t}\left(\Sigma\Lambda_{t}\right)+\frac{1}{2}(\left|\lambda_{t}\right|^{2}+\left|\Sigma\Lambda_{t}\right|^{2})+\Lambda_{t}^{\top}\Sigma\lambda_{t}\Big)dt$
$-\gamma J_{t}^{\alpha}e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}+J_{t}^{\alpha}\Lambda_{t}^{\top}dW_{t}=J_{t}^{\alpha}D_{t}(\alpha_{t})dt-\gamma J_{t}^{\alpha}e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}+J_{t}^{\alpha}\Lambda_{t}^{\top}dW_{t}.$
where the drift factor takes the form:
$D_{t}(\alpha)=$ $\frac{\gamma^{2}}{2}e^{\int_{t}^{T}2r(s)ds}\alpha^{\top}\alpha-\gamma e^{\int_{t}^{T}r(s)ds}\alpha^{\top}\big(\lambda_{t}+\Sigma\Lambda_{t}\big)+\frac{1}{2}\left|\lambda_{t}+\Sigma\Lambda_{t}\right|^{2}.$
Differentiating $D_{t}(\alpha)$ with respect to $\alpha$ and checking the second order condition, the maximizer $\alpha^{*}_{t}=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)$ for every $t\in[0,T]$ that is the strategy given by Equation (3.64). Evaluating the drift factor $D_{t}$ at $\alpha^{*}_{t}$ show that $D_{t}(\alpha^{*}_{t})$ vanishes to 0. Note $D_{t}(\alpha)$ is a quadratic function on $\alpha$ and $\gamma-1<0$. As $D_{t}(\alpha^{*}_{t})=0$, then $D_{t}(\alpha_{t})\geq 0$ for any admissible strategy $\alpha$. Moreover, solving the stochastic differential equation for $J_{t}^{\alpha}$ yields
$\hskip 85.35826pt\forall\,t\in[0,T],\quad J_{t}^{\alpha}=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)e^{\int^{t}_{0}D_{s}(\alpha_{s})ds}F_{t}^{\alpha}$
$\text{where}\quad F_{t}^{\alpha}=\mathcal{E}\Big(\int_{0}^{t}-\gamma e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)$
$\hskip 56.9055pt=\operatorname{exp}\Big(-\frac{1}{2}\int^{t}_{0}\Big(\gamma^{2}e^{2\int_{s}^{T}r(u)du}\left|\alpha_{s}\right|^{2}+\left|\Lambda_{s}\right|^{2}\Big)ds-\gamma\int_{0}^{t}e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big).$
Note $D_{t}(\alpha^{*}_{t})=0$, then $D_{t}(\alpha_{t})\geq 0$. By our assumptions on the admissible strategies 3.8 and Proposition 3.9, $\left(\alpha,\Lambda\right)\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ and thus, the stochastic exponential $F_{t}^{\alpha}$ is a local martingale (which follows from the basic properties of the Dooléans-Dade exponential). Therefore, there exists a sequence of stopping times $\{\tau_{n}\}_{n\geq 1}$ satisfying $\lim_{n\rightarrow\infty}\tau_{n}=T$, $\mathbb{P}$-a.s., such that that $F^{\alpha}_{t\wedge\tau_{n}}$ is a positive martingale for every $n$.
Furthermore, $\frac{\Gamma_{0}}{\gamma}\exp\big[-\gamma e^{\int^{T}_{0}r(u)du}x_{0}\big]e^{\int^{t}_{0}D_{s}(\alpha_{s})ds}$ is non-increasing. Therefore, $J^{\alpha}_{t\wedge\tau_{n}}$ is a supermartingale. Then for $s\leq t$, $\mathbb{E}[J^{\alpha}_{t\wedge\tau_{n}}|{\mathcal{F}}_{s}]\leq J^{\alpha}_{s\wedge\tau_{n}}$. It implies that for any set $A\in{\mathcal{F}}_{s}$,
$\mathbb{E}[J^{\alpha}_{t\wedge\tau_{n}}\mathds{1}_{A}]\leq\mathbb{E}[J^{\alpha}_{s\wedge\tau_{n}}\mathds{1}_{A}],\quad s\leq t.$
for every $n$. Since $\left(\exp\big[-\gamma e^{\int^{T}_{t\wedge\tau_{n}}r(u)du}X_{t\wedge\tau_{n}}\big]\right)_{n}$ is uniformly integrable (Definition 3.8) and $\Gamma$ is bounded, $\left(J^{\alpha}_{t\wedge\tau_{n}}\right)_{n}$ and $\left(J^{\alpha}_{s\wedge\tau_{n}}\right)_{n}$ are uniformly integrable. Let $n\rightarrow\infty$, then $\mathbb{E}[J^{\alpha}_{t}\mathds{1}_{A}]\leq\mathbb{E}[J^{\alpha}_{s}\mathds{1}_{A}]$. Then we deduce that $J^{\alpha}$ is a supermartingale for every arbitrary admissible strategy $\alpha$.
It remains to show that $J_{t}^{\alpha^{*}}$ is a true martingale for the optimal strategy $\alpha^{*}$ in which case $e^{\int_{0}^{t}D_{s}(\alpha_{s}^{*})ds}=1$ and hence $J_{t}^{\alpha^{*}}=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)F_{t}^{\alpha^{*}}$.
For $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a martingale by Lemma 5.1 with $g_{1,i}(t)=-\rho_{i}\theta_{i}+(1-\rho_{i}^{2})\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\in L^{\infty}([0,T],\mathbb{R})$ and $g_{2,i}(t)=-\sqrt{1-\rho_{i}^{2}}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\in L^{\infty}([0,T],\mathbb{R})$ for $1\leq i\leq d$. Subsequently, $J^{\alpha^{*}}_{t}$ is a true martingale. We have verified all conditions required by martingale optimality principle in Definition 3.2, except for the admissibility of $\alpha^{*}$, which follows directly from Step 1 in the proof of Theorem 3.12 below in the more general setting, as the arguments are similar. The proof is complete. $\Box$
Proof of Theorem 3.12:
Step 1 (Proof of Admissibility:) We begin by proving the admissibility of the candidate in equation (3.77) for the optimal portfolio strategy $\alpha^{*}$, and, on the way, we deduce the bound in equation (3.80).
In order to show that $\alpha^{*}$ is admissible, we have to show that (a), (b), (c) of Definition 3.8 hold. Part (b) is true because (3.53) has a unique solution (3.54) in terms of $(S,V,B)$. To prove (c), that is $\left\{\exp\big[-\gamma e^{\int^{T}_{\tau}r_{v}dv}X^{\alpha^{*}}_{\tau}\big]:\tau\text{ stopping time with values in }[0,T]\right\}$ is a uniformly integrable family, we only need to show
$\sup_{\tau}\mathbb{E}\Big[e^{-p\gamma e^{\int^{T}_{\tau}r(u)du}X^{\alpha^{*}}_{\tau}}\Big]<\infty,\text{ for some }p>1.$ (5.100)
Note that from (3.54), $X^{\alpha^{*}}_{\tau}=e^{\int^{\tau}_{0}r(s)ds}x_{0}+\int^{\tau}_{0}e^{\int^{\tau}_{s}r(u)du}\alpha_{s}^{*\top}\lambda_{s}\mathrm{d}s+\int^{\tau}_{0}e^{\int^{\tau}_{s}r(u)du}\alpha_{s}^{*\top}dB_{s},$ then:
$\mathbb{E}\Big[e^{-p\gamma e^{\int^{T}_{\tau}r(u)du}X^{\alpha^{*}}_{\tau}}\Big]$ $\leq e^{-p\gamma e^{\int^{T}_{0}r(u)du}x_{0}}\mathbb{E}\Big[\exp\Big(-p\gamma\int^{\tau}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{*\top}\lambda_{s}ds-p\gamma\int^{\tau}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{*\top}dB_{s}\Big)\Big]$
$\leq e^{-p\gamma e^{\int^{T}_{0}r(u)du}x_{0}}\mathbb{E}\Big[\exp\Big(p\gamma\int^{T}_{0}e^{\int^{T}_{s}r(u)du}\big|\alpha_{s}^{*\top}\lambda_{s}\big|ds+p^{2}\gamma^{2}\int^{T}_{0}e^{\int^{T}_{s}2r(u)du}\left|\alpha_{s}^{*}\right|^{2}ds\Big)$
$\hskip 42.67912pt\times\exp\Big(-p^{2}\gamma^{2}\int^{\tau}_{0}e^{\int^{T}_{s}2r(u)du}\left|\alpha_{s}^{*}\right|^{2}ds-p\gamma\int^{\tau}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{*\top}dB_{s}\Big)\Big]$
so that by Hölder’s inequality, followed by the elementary inequality $ab\leq(a^{2}+b^{2})/2$ in the third line, we may write
$\,\sup_{\tau}\mathbb{E}\Big[e^{-p\gamma e^{\int^{T}_{\tau}r(u)du}X^{\alpha^{*}}_{\tau}}\Big]\leq e^{-p\gamma e^{\int^{T}_{0}r(u)du}x_{0}}\mathbb{E}\Big[\exp\Big(\int^{T}_{0}(2p^{2}\gamma^{2}e^{\int^{T}_{s}2r(u)du}\left|\alpha_{s}^{*}\right|^{2}+2p\gamma e^{\int^{T}_{s}r(u)du}\big|\alpha_{s}^{*\top}\lambda_{s}\big|)ds\Big)\Big]^{\frac{1}{2}}$
$\hskip 128.0374pt\times\sup_{\tau}\mathbb{E}\Big[\exp\Big(-2p^{2}\gamma^{2}\int^{\tau}_{0}e^{\int^{T}_{s}2r(u)du}\left|\alpha_{s}^{*}\right|^{2}ds-2p\gamma\int^{\tau}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{*\top}dB_{s}\Big)\Big]^{\frac{1}{2}}$
$\;\leq\frac{e^{-p\gamma e^{\int^{T}_{0}r(u)du}x_{0}}}{\sqrt{2}}\Big(\mathbb{E}\big[\exp\big(4p^{2}\gamma^{2}\int^{T}_{0}e^{\int^{T}_{s}2r(u)du}\left|\alpha_{s}^{*}\right|^{2}ds\big)\big]+\mathbb{E}\big[\exp\big(4p\gamma\int^{T}_{0}e^{\int^{T}_{s}r(u)du}\big|\alpha_{s}^{*\top}\lambda_{s}\big|ds\big)\big]\Big)^{\frac{1}{2}}\times 1$
which is finite since on the first hand, the Doléans-Dade exponential $\mathcal{E}\big(-\int_{0}^{\cdot}2p\gamma e^{\int^{T}_{s}r(u)du}\alpha_{s}^{*\top}dB_{s}\big)$ satisfies the Novikov’s condition by virtue of (3.27), and is therefore a true martingale (see also Lemma 5.2) with Expectation $1$. Then by optional sampling theorem with the fact that $\tau\leq T$, it follows that the supremum is bounded by $1$.
On the others hand, we used Jensen’s inequality to obtain the bound
$e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}^{*}\right|^{2}=\frac{1}{\gamma^{2}}|\lambda_{s}+\Sigma\Lambda_{s}|^{2}\leq\frac{2}{\gamma^{2}}(|\lambda_{s}|^{2}+|\Sigma\Lambda_{s}|^{2})\leq\frac{2}{\gamma^{2}}(1+|\Sigma|^{2})(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2})$
so that the first term in the bracket is finite thanks to condition (3.27) with constant $a(p)=2(8p^{2}{-2p})\left(1+|{\Sigma}|^{2}\right)$ after noticing that $8p^{2}\leq 2(8p^{2}{-2p})$ for all $p>1$.
$\mathbb{E}\Big[\exp\big(4p^{2}\gamma^{2}\int^{T}_{0}e^{\int^{T}_{s}2r(u)du}\left|\alpha_{s}^{*}\right|^{2}ds\big)\Big]\leq\mathbb{E}\left[\exp\left({a(p)\int_{0}^{T}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds}\right)\right]<\infty,$ (5.101)
Still owing to the elementary inequality $|ab|\leq(|a|^{2}+|b|^{2})/2$, we have
$e^{\int^{T}_{s}r(u)du}|\alpha_{s}^{*\top}\lambda_{s}|=\frac{1}{\gamma}\big(|\lambda_{s}|^{2}+|(\Sigma\Lambda_{s})^{\top}\lambda_{s}|\big)\leq\frac{1}{\gamma}(|\lambda_{s}|^{2}+|\Sigma|\frac{|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}}{2})\leq\frac{1}{2\gamma}(2+|\Sigma|)(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2})$
which ensures that the second term in the bracket is finite i.e.
$\mathbb{E}\Big[\exp\big(4p\gamma\int^{T}_{0}e^{\int^{T}_{s}r(u)du}\big|\alpha_{s}^{*\top}\lambda_{s}\big|ds\big)\Big]\leq\mathbb{E}\left[\exp\left({a(2p)\int_{0}^{T}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds}\right)\right]<\infty,$ (5.102)
thanks to condition (3.27) with constant $a(2p)=2p\left(2+|{\Sigma}|\right)$. We are left to prove that $\alpha^{*}\in L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. We have:
$\mathbb{E}\left[\int_{0}^{T}|\alpha_{s}^{*}|^{2}ds\right]$ ${=}\mathbb{E}\left[\int_{0}^{T}\frac{1}{\gamma^{2}}e^{-2\int_{s}^{T}r(s)ds}|\lambda_{s}+\Sigma\Lambda_{s}|^{2}ds\right]\leq\mathbb{E}\left[\frac{2}{\gamma^{2}}(1+|\Sigma|^{2})\int_{0}^{T}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds\right]<\infty,$
where the last term is finite due to condition (3.27) and the inequality $|z|^{q}\leq c_{q}e^{|z|},\;\forall\,q\geq 1$. It becomes straightforward that $\alpha^{*}$ in equation (3.77) is admissible.
Step 2 (Proof of the equality (1):) Let’s us consider the Problem (3.56) with an arbitrary strategy $\alpha\in\mathcal{A}$. To ease notations, we set $Y_{t}:=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\right)$ and $h_{t}=\lambda_{t}+\Sigma\Lambda_{t}$. For any admissible strategy $\alpha\in\mathcal{A}$, Itô’s lemma combined with the property of $\Gamma$ in Equation (3.71) of Proposition 3.11 yield:
$d\big(\Gamma_{t}Y_{t}\big)\;$ $=\Gamma_{t}Y_{t}\Big(D_{t}(\alpha_{t})dt-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}+\Lambda_{t}^{\top}dW_{t}\Big).$
where the drift factor takes the form $D_{t}(\alpha_{t})=\frac{\gamma^{2}}{2}e^{\int_{t}^{T}2r(s)ds}\alpha_{t}^{\top}\alpha_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}h_{t}+\frac{1}{2}h_{t}^{\top}h_{t}.$ As a consequence, using $\Gamma_{T}=1$, we get
$\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)e^{\int^{T}_{0}D_{s}(\alpha_{s})ds}F_{T}^{\alpha}.$ (5.103)
where $F_{T}^{\alpha}$ is an $(\mathbb{F},\mathbb{P})$-local martingale as $(\alpha,\Lambda)$ are in $L^{2,{loc}}_{\mathbb{F}}([0,T])$ and write
$F_{T}^{\alpha}$ $=\mathcal{E}\Big(\int_{0}^{T}-\gamma e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)$
$=\operatorname{exp}\Big(-\frac{1}{2}\int^{T}_{0}\Big(\gamma^{2}e^{2\int_{s}^{T}r(u)du}\left|\alpha_{s}\right|^{2}+\left|\Lambda_{s}\right|^{2}\Big)ds-\gamma\int_{0}^{T}e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\int_{0}^{T}\Lambda_{s}^{\top}dW_{s}\Big).$
Now, for $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a $\mathbb{P}$-martingale with expectation $1$ by Lemma 5.1 with $g_{1,i}(t)=-\rho_{i}\theta_{i}+(1-\rho_{i}^{2})\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\in L^{\infty}([0,T],\mathbb{R})$ and $g_{2,i}(t)=-\sqrt{1-\rho_{i}^{2}}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\in L^{\infty}([0,T],\mathbb{R})$ for $1\leq i\leq d$. Consequently, inserting the candidate (3.77) for the optimal strategy $\alpha^{*}$ in the above Equation (5.103), and observing that $D_{s}(\alpha_{s}^{*})=0\;\forall\,s\in[0,T]$, then taking the expectation, its boils down that:
$\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha^{*}}_{T}\right)\big]}=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}{\big[F_{T}^{\alpha^{*}}\big]}=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right).$
where the last inequality comes from the fact that $F_{T}^{\alpha^{*}}$ is a true martingale. This completes the first part $(1)$ of the proof. It remains to show the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$.
Step 3 (Proof of the inequality (2):) Note that the SDE for the wealth process (3.54) can be solved explicitly so as to admit the representation
$X_{T}^{\alpha}=e^{\int^{T}_{0}r(s)ds}x_{0}+\int^{T}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}\mathrm{d}s+\int^{T}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}dB_{s}.$ (5.104)
Since $\alpha_{t}$ is bounded by assumption and $s\to e^{\int^{T}_{s}r(u)du}\in L^{\infty}([0,T],\mathbb{R})$, we can define a new probability measure $\mathbb{Q}$ with Radon-Nikodym density
${Z}_{T}:=\frac{d{\mathbb{Q}}}{d\mathbb{P}}|_{\mathcal{F}_{T}}=\exp\Big(-\gamma\int_{0}^{T}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\,dB_{s}-\frac{\gamma^{2}}{2}\int_{0}^{T}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}ds\Big),$ (5.105)
by Lemma 5.2. Recalling that $U(X_{T}^{\alpha})=-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)$, by Girsanov’s theorem, we may write:
$\mathbb{E}_{x_{0},V_{0}}\Big[U(X_{T}^{\alpha})\Big]$ $\,=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}\big(-\gamma e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}\big)ds\Big)\Big],$
$=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}^{\alpha}ds\Big)\Big],$
with the real-valued deterministic process $F_{s}$ given by
$\forall\,s\in[0,T],\quad F_{s}^{\alpha}$ $=-\gamma e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}$
$=F_{s}^{\alpha^{*}}-\gamma e^{\int^{T}_{s}r(u)du}\hat{\alpha}_{s}^{\top}\lambda_{s}+\gamma^{2}e^{2\int^{T}_{s}r(u)du}\alpha_{s}^{*\top}\hat{\alpha}_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\hat{\alpha}_{s}\right|^{2}$
where we write the arbitrary admissible strategy $(\alpha_{t})_{t\in[0,T]}$ in terms of the optimal strategy $\alpha^{*}_{t}=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)$ for every $t\in[0,T]$ and some remainder $(\hat{\alpha}_{t})_{t\in[0,T]}$ i.e. $\alpha_{t}=\alpha_{t}^{*}+\hat{\alpha}_{t}$ for every $t\in[0,T]$. Note that, on the first hand, substituting the optimal strategy $\alpha^{*}$ in (3.77) yields $\forall\,s\in[0,T]$
$F_{s}^{\alpha^{*}}=-\left(\lambda_{s}+\Sigma\Lambda_{s}\right)^{\top}\lambda_{s}+\frac{1}{2}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}=\frac{1}{2}(-\left|\lambda_{s}\right|^{2}+\left|\Sigma\Lambda_{s}\right|^{2})=\sum_{i=1}^{d}\Big(-\frac{\theta_{i}^{2}}{2}+\frac{1}{2}\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}$
and on the other hand, recalling that $\alpha_{s}:=\sqrt{\operatorname{diag}(V_{s})}^{\top}\pi_{s}$ and defining $\hat{\pi}:=\pi-\pi^{*}$ so that $\pi_{s}=\pi_{s}^{*}+\hat{\pi}_{s}$, with $\pi_{s}^{*}=\frac{1}{\gamma}e^{-\int_{s}^{T}r(u)du}\left(\theta+\Sigma\nu\varsigma(s)\psi(T-s)\right)$ so that $\pi_{s,i}^{*}=\frac{1}{\gamma}e^{-\int_{s}^{T}r(u)du}(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s))$ for $i=1,\ldots,d$ and consequently, $\forall\,s\in[0,T],$
$F_{s}^{\alpha}$ $=F_{s}^{\alpha^{*}}+\sum_{i=1}^{d}\Big(-\gamma e^{\int^{T}_{s}r(u)du}\theta_{i}\hat{\pi}_{s,i}+\gamma^{2}e^{2\int^{T}_{s}r(u)du}\pi_{s,i}^{*}\hat{\pi}_{s,i}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
$=\sum_{i=1}^{d}\Big(-\frac{\theta_{i}^{2}}{2}+\frac{1}{2}\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}+\gamma e^{\int^{T}_{s}r(u)du}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
Using Equation (3.66) reading $\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))=(\psi^{i}\star r^{i})^{\prime}(T-s)$ owing to Lemma 5.3 $(a)$ for $i=1,\ldots,d$, where $r^{i}$ the resolvent of the first kind of the integral kernel $K_{i}$, one gets:
$\forall\,s\in[0,T],\;F_{s}$ $=\sum_{i=1}^{d}(\psi^{i}\star r^{i})^{\prime}(T-s)V^{i}_{s}-\Big(F_{i}(s,\psi(T-s))-\frac{1}{2}\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}$ (5.106)
$\hskip 42.67912pt+\sum_{i=1}^{d}\Big(\gamma e^{\int^{T}_{s}r(u)du}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
Under the probability measure $\mathbb{Q}$ defined in (5.105), the processes $\widehat{B}_{t}:=B_{t}+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\alpha_{s}ds$ and
$\widehat{W}_{t}:=\Sigma\widehat{B}_{t}+\sqrt{I-\Sigma^{\top}\Sigma}B^{\perp}_{t}=W_{t}+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\Sigma\alpha_{s}ds=W_{t}+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\Sigma\alpha^{*}_{s}ds+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\Sigma\hat{\alpha}_{s}ds,$
are standard Wiener processes thanks to the Cameron–Martin–Girsanov and Girsanov theorem respectively. The dynamics of the variance process $V$ under $\mathbb{Q}$ can thus be written as
$\,V_{t}^{i}=g_{0}^{i}(t)+\int_{0}^{t}K_{i}(t-s)H_{s}^{i}ds+\int_{0}^{t}K_{i}(t-s)\nu_{i}\varsigma^{i}(s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s},\quad i=1,\ldots,d;\quad\text{with}$
$H_{s}^{i}:=(DV_{s})_{i}-\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)e^{\int^{T}_{s}r(u)du}(\pi_{s,i}^{*}+\hat{\pi}_{s,i})V_{s}^{i},\quad i=1,\ldots,d.$
$\hskip 17.07182pt=-\rho_{i}\theta_{i}\nu_{i}\varsigma^{i}(s)V_{s}^{i}+(DV_{s})_{i}-(\rho_{i}\nu_{i}\varsigma^{i}(s))^{2}\psi^{i}(T-s)V_{s}^{i}-\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)e^{\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}V_{s}^{i}$
Now, we insert the dynamics of $V^{i}$ into the expression $\int_{t}^{T}(\psi^{i}\star r^{i})^{\prime}(T-s)V^{i}_{s}ds$ and simplify using Lemma 5.3 $(b)$, where we reads $H^{i}$ and the stochastic term pathwise (i.e. $\omega$ by $\omega$) as for example $g(s)=H^{i}_{s}(\omega)$. Its boils down that:
$\,\int_{t}^{T}\sum_{i=1}^{d}(\psi^{i}\star r^{i})^{\prime}(T-s)V^{i}_{s}ds=\sum_{i=1}^{d}\int_{t}^{T}(\psi^{i}\star r^{i})^{\prime}(T-s)g_{0}^{i}(s)ds+\int_{t}^{T}\psi^{i}(T-s)\big(H_{s}^{i}ds+\nu_{i}\varsigma^{i}(s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\big)$
$\hskip 42.67912pt=\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{1}{2}\theta_{i}^{2}+F_{i}(s,\psi(T-s))\big)g_{0}^{i}(s)ds+\sum_{i=1}^{d}\int_{t}^{T}\psi^{i}(T-s)\big(H_{s}^{i}ds+\nu_{i}\varsigma^{i}(s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\big).$
Noting by a change of variables that
$\sum_{j=1}^{d}\psi^{j}(T-s)(DV_{s})_{j}=\sum_{j=1}^{d}\psi^{j}(T-s)\sum_{i=1}^{d}D_{ji}V_{s}^{i}=\sum_{i=1}^{d}V_{s}^{i}\sum_{j=1}^{d}D_{ji}\psi^{j}(T-s)=\sum_{i=1}^{d}(D^{\top}\psi)_{i}(T-s)V_{s}^{i},$
one gets using Equation (3.69)
$\sum_{i=1}^{d}\int_{0}^{T}\psi^{i}(T-s)H_{s}^{i}ds$ $=\sum_{i=1}^{d}\int_{0}^{T}\Big(F_{i}(s,\psi^{i}(T-s))-\frac{\nu_{i}^{2}}{2}\big(1+\rho_{i}^{2}\big)(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V_{s}^{i}\,ds$
$-\gamma\sum_{i=1}^{d}\int_{0}^{T}e^{\int^{T}_{s}r(u)du}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}V_{s}^{i}ds.$
Replacing back all in (5.106), some terms cancel out and we end up with
$\int_{0}^{T}F_{s}\,ds$ $=\sum_{i=1}^{d}\int_{0}^{T}\Big(-\frac{1}{2}\theta_{i}^{2}+F_{i}(s,\psi(T-s))\Big)g_{0}^{i}(s)\,ds+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds$
$\qquad+\sum_{i=1}^{d}\int_{0}^{T}\Big(-\frac{\nu_{i}^{2}}{2}\big(\varsigma^{i}(s)\psi^{i}(T-s)\big)^{2}V_{s}^{i}\,ds+\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}\,d\widehat{W}^{i}_{s}\Big).$
Hence we obtain
$\,\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)\Big]=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}^{\alpha}ds\Big)\Big],$
$\hskip 128.0374pt=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 79.6678pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}$
$\hskip 170.71652pt+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds\Big)\Big].$ (5.107)
Remark: (Alternative Proof of the equality (1):) At this stage, note that if the admissible strategy $\alpha$ is optimal, i.e. $\alpha=\alpha^{*}$ so that $\hat{\pi}\equiv 0$ and $\mathbb{Q}\equiv\mathbb{Q}^{*}$ in (5.105), then Equation (5.107) becomes:
$\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha^{*}}_{T}\right)\Big]=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 62.59596pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}^{*}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{*i}_{s}\Big)\Big].$
Since $\psi$ is continuous, it is bounded and therefore the stochastic exponential is a true $\mathbb{Q}^{*}$- martingale with expectation $1$ by Lemma 5.1 with $g_{2}=0$ and $g_{1,i}(s)=\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)$ . Thus we get
$\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha^{*}}_{T}\right)\Big]=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$
This completes the first part of the proof (equality (1)).
Back to the proof of the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$, still from Equation (5.107), since $V_{s}$ is positive definite, the term $\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds$ or equivalently $\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\alpha}_{s,i}^{2}\,ds$ has to be more than or equal to 0 so that
$\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)\Big]\leq-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 62.59596pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\Big)\Big].$
Since the stochastic exponential is a $\mathbb{Q}$-martingale with expectation $1$ still by Lemma 5.1, we finally obtain that for every $\alpha\in\mathcal{A}$
$\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)\big]}\leq-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$
which completes the proof and we are done. $\Box$
Acknowledgement: I thank Gilles Pagès, Mathieu Rosenbaum and Dro Sigui for insightful discussions.